### v1

# Compatibility of Drinfeld presentations for split affine Kac–Moody quantum symmetric pairs

Abstract.
Let $(\mathbf{U},\mathbf{U}^{\imath})$ be a split affine quantum symmetric pair of type $\mathsf{B}_{n}^{(1)},\mathsf{C}_{n}^{(1)}$ or $\mathsf{D}_{n}^{(1)}$. We prove factorization and coproduct formulae for the Drinfeld–Cartan operators $\Theta_{i}(z)$ in the Lu–Wang Drinfeld-type presentation, generalizing the type $\mathsf{A}_{n}^{(1)}$ result from [Prz23]. As an application, we show that a boundary analogue of the $q$-character map, defined via the spectra of these operators, is compatible with the usual $q$-character map. As an auxiliary result, we also produce explicit reduced expressions for the fundamental weights in the extended affine Weyl groups of classical types.
Key words and phrases:
Quantum symmetric pairs, coideal subalgebras, $\imath$quantum groups, Kac–Moody algebras, $q$-Onsager algebra, Drinfeld presentation, Drinfeld polynomials, braid group action
2020 Mathematics Subject Classification:
17B37, 17B67, 81R10
The first author was supported by the Austrian Science Fund (FWF): P-34602, Grant DOI: 10.55776/P34602, and PAT 9039323, Grant-DOI 10.55776/PAT9039323. The second author was supported by the EPSRC grant No. EP/W022834/1 Kac–Moody quantum symmetric pairs, KLR algebras and generalized Schur–Weyl duality.
1. Introduction
Quantum affine algebras admit three distinct presentations: the ‘Drinfeld–Jimbo’, ‘new Drinfeld’ and ‘RTT’ realizations. The first can be seen as a quantization of the usual Serre-type presentation of a Kac-Moody Lie algebra $\widehat{{\mathfrak{g}}}$, while the second is a quantization of the central extension presentation of $\widehat{{\mathfrak{g}}}$. The interplay between these different realizations was studied and described precisely in [DF93, FM02]. One of the most interesting features of the new Drinfeld presentation is that it exhibits a large, infinitely generated, commutative subalgebra of $U_{q}(L{\mathfrak{g}})$. The spectra of the generators of this subalgebra, often called Drinfeld–Cartan operators $\phi^{\pm}_{i,m}$, play a key role in the classification of finite dimensional representation of $U_{q}(L{\mathfrak{g}})$ via Drinfeld polynomials [CP91, CP94], and $q$-character theory [FR99, FM01].
Recently, Lu and Wang [LW21], building on the work of Baseilhac and Kolb [BK20] in rank one (see also [Zha22, LWZ24]), have constructed a Drinfeld-type presentation for split and certain quasi-split affine quantum symmetric pair coideal subalgebras. The Lu–Wang presentation also exhibits a large commutative subalgebra, generated by $\Theta_{i,m}$. This very significant development raises many questions. Here we focus on the following three. Firstly, what is the relationship between the generators $\Theta_{i,m}$ and the usual Drinfeld–Cartan operators $\phi^{\pm}_{i,m}$? Secondly, how to describe the coproduct of $\Theta_{i,m}$? Thirdly, can the Lu–Wang presentation be used to define a ‘boundary’ analogue of $q$-characters?
The first two questions were answered by the second author [Prz23] for split affine quantum symmetric pairs of type $\mathsf{A}_{n}^{(1)}$. In particular, it was shown that the generating series $\boldsymbol{\grave{\Theta}}_{i}(z)$ have the following factorization property:
(1.1) $\boldsymbol{\grave{\Theta}}_{i}(z)\equiv\boldsymbol{\phi}_{i}^{-}(z^{-1})\boldsymbol{\phi}_{i}^{+}(Cz)\quad\mod{U_{q}(L{\mathfrak{g}})_{+}}[\negthinspace[z]\negthinspace],$
where $U_{q}(L{\mathfrak{g}})_{+}$ is the Drinfeld positive half of $U_{q}(L{\mathfrak{g}})$. Moreover, the series $\boldsymbol{\grave{\Theta}}_{i}(z)$ are approximately group-like, in the sense that
(1.2) $\Delta(\boldsymbol{\grave{\Theta}}_{i}(z))\equiv\boldsymbol{\grave{\Theta}}_{i}(z)\otimes\boldsymbol{\grave{\Theta}}_{i}(z)\quad\mod{U_{q}(L{\mathfrak{g}})\otimes U_{q}(L{\mathfrak{g}})_{+}}[\negthinspace[z]\negthinspace].$
The main result of the present paper is a generalization of the two results above to split affine quantum symmetric pairs of types $\mathsf{B}_{n}^{(1)},\mathsf{C}_{n}^{(1)}$ and $\mathsf{D}_{n}^{(1)}$. As an application, we also propose a definition of boundary $q$-characters and prove that they are, in a suitable sense, compatible with the usual $q$-characters.
1.1. Proof strategy
Our proof is based on an algebraic and computational approach, using the Drinfeld–Jimbo and new Drinfeld presentations only. In particular, we do not use any information available through the RTT presentation. Since the overall proof is quite technical, let us summarize the main steps below.
Step 1: compute explicit reduced expressions for the fundamental weights $\omega_{i}$. This is carried out in §3. Whenever $\alpha_{i}$ has multiplicity one in the highest root, the expressions can be derived from a formula in [Bou02]. In the remaining cases, we consider the extended affine Weyl group as a group of affine transformations of ${\mathbb{R}}^{n}$ and compute that our expressions act as the correct affine shifts. To show that they are reduced, we calculate their lengths and compare them with a formula from [IM65].
Step 2: express $\mathbf{T}_{\omega^{\prime}_{i}}(B_{i})$ as an explicit polynomial in the variables $B_{j}$. As in the usual Drinfeld presentation, the generators in the Lu–Wang presentation are constructed using a braid group action (which differs from the ordinary Lusztig action). More precisely, they are constructed by repeatedly applying the braid group operators $\mathbf{T}_{\omega_{i}}$ to the usual Kolb–Letzter generators of the coideal subalgebra. Using the explicit formulae for $\omega_{i}$ from Step $1$, we are able to express $\mathbf{T}_{\omega^{\prime}_{i}}(B_{i})$ by means of certain recursively defined polynomials in the variables $B_{j}$. This is carried out in §7, based on the auxiliary calculations from §4.
Step 3: show that $\mathbf{T}_{\omega^{\prime}_{i}}(B_{i})\equiv T_{\omega^{\prime}_{i}}(B_{i})$. More precisely, we show that the Lu–Wang braid group operator $\mathbf{T}_{\omega^{\prime}_{i}}$ coincides with the usual Lusztig braid group operator $T_{\omega^{\prime}_{i}}$ on $B_{i}$, modulo a certain subalgebra of $U_{q}(L{\mathfrak{g}})_{+}$. This is achieved in §7 (Theorem 7.1) in the following way. Recall that $B_{j}$ is a linear combination of $F_{j}$ and $E_{j}K_{j}^{-1}$. Writing $\mathbf{T}_{\omega^{\prime}_{i}}(B_{i})$ as a polynomial as in Step $2$, and substituting for each occurrence of $B_{j}$ either $F_{j}$ or $E_{j}K_{j}^{-1}$, we can express $\mathbf{T}_{\omega^{\prime}_{i}}(B_{i})$ as a sum of homogeneous terms in the Drinfeld gradation. The main problem is to show that, except for $T_{\omega^{\prime}_{i}}(E_{i}K_{i}^{-1})$, the terms which are not in $U_{q}(L{\mathfrak{g}})_{+}$ vanish. This is solved using the theory of good polynomials developed in §5, with the aid of auxiliary calculations from §6.
Step 4: deduce the factorization and coproduct formulae. The main idea is that, modulo $U_{q}(L{\mathfrak{g}})_{+}$, we can reduce the problem to rank one subalgebras, and apply the results from [Prz23]. We show that, to perform this reduction, we only need to verify a single commutation condition. This criterion is then checked explicitly in type $\mathsf{D}_{n}^{(1)}$. The same methods carry over to types $\mathsf{B}_{n}^{(1)}$ and $\mathsf{C}_{n}^{(1)}$. This is achieved in §8.
1.2. Reduced expressions for fundamental weights
In Step $1$ above, we compute explicit reduced expressions for the fundamental weights in the extended affine Weyl groups of classical types. Somewhat surprisingly, to our knowledge, such explicit expressions have not appeared in the literature before. As this result is interesting in its own right, we collect the expressions in a list below for quick reference, see Table 1.
Type Fundamental weight $\omega_{i}$ Range of $i$
$\mathsf{A}_{n}^{(1)}$ $\pi^{i}[n{\shortminus}i{+}1,n]\cdots[2,i{+}1][1,i]$ $1\leq i\leq n$
$\mathsf{B}_{n}^{(1)}$ $\pi_{1}[1,n]\big{(}s_{0}[2,n][1,n]\big{)}^{\frac{i{\shortminus}1}{2}}[{n{\shortminus}i},{n{\shortminus}1}]\cdots[1,i]$ $1\leq i\leq n$, $i$ is odd
$\big{(}s_{0}[2,n][1,n]\big{)}^{\frac{i}{2}}[{n{\shortminus}i},{n{\shortminus}1}]\cdots[1,i]$ $2\leq i\leq n$, $i$ is even
$\mathsf{C}_{n}^{(1)}$ $\pi_{n}s_{n}[n{\shortminus}1,n]\cdots[1,n]$ $i=n$
$(s_{0}[1,n])^{i}[n{\shortminus}i,n{\shortminus}1]\cdots[2,i{+}1][1,i]$ $i<n$
$\mathsf{D}_{n}^{(1)}$ $\big{(}s_{0}[2,n{\shortminus}1][1,n{\shortminus}2]s_{n}\big{)}^{\frac{i}{2}}[{n{\shortminus}i},{n{\shortminus}1}]\cdots[1,i]$ $22\leq i\leq n{\shortminus}$ 2, $i$ is even
$\pi_{n{\shortminus}1}r_{n{\shortminus}2}r_{n{\shortminus}4}\cdots r_{3}s_{n}[1,{n{\shortminus}1}]$ $i=n{\shortminus}1$, $i$ is even
$\pi_{n}r_{n{\shortminus}2}r_{n{\shortminus}4}\cdots r_{2}s_{n}$ $i=n$, $i$ is even
$\pi_{1}[1,n{\shortminus}2]s_{n}\big{(}s_{0}[2,n{\shortminus}1][1,n{\shortminus}2]s_{n}\big{)}^{\frac{i{\shortminus}1}{2}}[{n{\shortminus}i},{n{\shortminus}1}]\cdots[1,i]$ $21\leq i\leq n{\shortminus}$ 2, $i$ is odd
$\pi_{n{\shortminus}1}s_{n{\shortminus}1}s_{n{\shortminus}2}r_{n{\shortminus}3}r_{n{\shortminus}5}\cdots r_{3}s_{n}[1,{n{\shortminus}1}]$ $i=n{\shortminus}1$, $i$ is odd
$\pi_{n}s_{n{\shortminus}1}s_{n{\shortminus}2}r_{n{\shortminus}3}r_{n{\shortminus}5}\cdots r_{2}s_{n}$ $i=n$, $i$ is odd
Table 1. Reduced expressions for fundamental weights $\omega_{i}$ in extended affine Weyl groups of classical type. Here $[k,l]=s_{k}s_{k+1}\cdots s_{l}$ and $r_{m}=s_{n}[m,n{\shortminus}1][m{\shortminus}1,n{\shortminus}2]$.
1.3. Application to $q$-characters
The notion of $q$-characters can be defined in at least three equivalent ways: via the universal $R$-matrix, via the spectrum of Drinfeld–Cartan operators, or via Nakajima’s quiver varieties. The quantum symmetric pair analogues of these approaches are objects of intense study, see, e.g., [AV22a, AV22b, AV24] for the latest developments on the universal $K$-matrix, and [Li19] for a quiver variety approach to symmetric pairs. Nevertheless, it appears that the geometric and integrable systems methods are not yet advanced enough to yield a satisfactory theory of $q$-characters for quantum symmetric pairs. The main obstacle in the latter case is the unavailability of a Khoroshkin-Tolstoy-Levendorsky-Soibelman-Stukopkin-Damiani-type factorization of the (affine) universal $K$-matrix.
In light of the aforementioned difficulties, we propose to define boundary $q$-characters directly via the Lu–Wang presentation instead. More precisely, we consider the generalized eigenspace decomposition of a finite dimensional representation with respect to the action of the operators $\Theta_{i,m}$, and let boundary $q$-characters encode the multiplicities of such eigenspaces. This is equivalent to taking the trace of a certain operator in the completion ${\mathbf{U}^{\imath}\otimes U_{q}(\widetilde{\mathfrak{h}})}[\negthinspace[z]\negthinspace]$, yielding a map $\chi_{q}^{\imath}\colon\operatorname{Rep}\mathbf{U}^{\imath}\to{U_{q}(\widetilde{\mathfrak{h}})}[\negthinspace[z]\negthinspace]$. We apply (1.1)–(1.2) to show that $\chi_{q}^{\imath}$ is compatible with the usual $q$-character map (Corollary 9.3). More precisely, we show that the diagram
[…]
commutes if ${U_{q}(\widetilde{\mathfrak{h}})}[\negthinspace[z]\negthinspace]$ is endowed with an appropriate ‘twisting’ action of ${\mathbb{Z}}[Y_{i,a}^{\pm 1}]_{i\in I,a\in{\mathbb{C}}^{\times}}$. In particular, this result yields an easy way to compute the boundary $q$-characters of restriction representations.
Acknowledgements
We would like to thank Andrea Appel, Pascal Baseilhac and Bart Vlaar for insightful discussions and comments.
2. Preliminaries
We work over the field of complex numbers and assume that $q\in{\mathbb{C}}^{\times}$ is not a root of unity throughout.
2.1. Quantum affine algebras
Let $\mathbb{I}_{0}=\{1,\cdots,n\}$ and $\mathbb{I}=\mathbb{I}_{0}\cup\{0\}$. Let ${\mathfrak{g}}$ be a simple Lie algebra with Cartan matrix $(a_{ij})_{i,j\in\mathbb{I}_{0}}$, and $\widehat{{\mathfrak{g}}}$ the corresponding untwisted affine Lie algebra with affine Cartan matrix $(a_{ij})_{i,j\in\mathbb{I}}$. Let $d_{i}$ be relatively prime positive integers such that $(d_{i}a_{ji})_{i,j\in\mathbb{I}_{0}}$ is a symmetric matrix, and set $q_{i}=q^{d_{i}}$. We use Bourbaki’s conventions [Bou02] for the Cartan matrix in non-simply laced types. i.e.,
$\left(\begin{array}[]{rrrrrrr}2&-1&\ldots&0&0\\ -1&2&\ldots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\ldots&2&-2\\ 0&0&\ldots&-1&2\end{array}\right)\quad,\quad\left(\begin{array}[]{rrrrrrr}2&-1&\ldots&0&0\\ -1&2&\ldots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\ldots&2&-1\\ 0&0&\ldots&-2&2\end{array}\right)$
are the Cartan matrices of types $\mathsf{B}_{n}$ and $\mathsf{C}_{n}$, respectively.
We use standard notations and conventions regarding root systems, weight lattices, Weyl groups, etc., as in, e.g., [LW21, §3.1]. In particular, we let $\alpha_{i}$ ($i\in\mathbb{I}$) denote the simple roots of $\widehat{{\mathfrak{g}}}$; let $\theta$ denote the highest root, and $\delta$ the basic imaginary root; let $P$ and $Q$ denote the weight and root lattices of ${\mathfrak{g}}$, respectively; and let $\omega_{i}\in P$ ($i\in\mathbb{I}_{0}$) be the fundamental weights of ${\mathfrak{g}}$. Note that, according to Bourbaki’s conventions, $s_{\alpha_{i}}(\alpha_{j})=\alpha_{j}-a_{ji}\alpha_{i}$.
The quantum affine algebra $U_{q}(\widehat{\mathfrak{g}})$ is the algebra with generators $e_{i}^{\pm},K_{i}^{\pm 1}$ $(i\in\mathbb{I})$ and relations:
(2.1) $K_{i}K_{i}^{-1}$ $=K_{i}^{-1}K_{i}=1,$
(2.2) $K_{i}K_{j}$ $=K_{j}K_{i},$
(2.3) $K_{i}e_{j}^{\pm}$ $=q_{i}^{\pm a_{ji}}e_{j}^{\pm}K_{i},$
(2.4) $[e_{i}^{+},e_{j}^{-}]$ $=\delta_{ij}\frac{K_{i}-K_{i}^{-1}}{q_{i}-q_{i}^{-1}},$
(2.5) $\mathsf{Serre}_{ij}(e_{i}^{\pm},e_{j}^{\pm})$ $=0\qquad(i\neq j),$
where
[…]
We will also sometimes abbreviate
$E_{i}=e_{i}^{+},\qquad F_{i}=e_{i}^{-}.$
We use the standard notation for divided powers, i.e., $(e^{\pm}_{i})^{(r)}=(e^{\pm}_{i})^{r}/[r]_{q_{i}}!$.
The algebra $U_{q}(\widehat{\mathfrak{g}})$ is a Hopf algebra, with the coproduct
(2.6) $\Delta(e_{i}^{+})=e_{i}^{+}\otimes 1+K_{i}\otimes e_{i},\quad\Delta(e_{i}^{-})=e_{i}^{-}\otimes K_{i}^{-1}+1\otimes e_{i}^{-},\quad\Delta(K_{i}^{\pm 1})=K_{i}^{\pm 1}\otimes K_{i}^{\pm 1}.$
The counit is given by
$\varepsilon(e_{i}^{\pm})=0,\quad\varepsilon(K_{i}^{\pm 1})=1.$
The quantum loop algebra $U_{q}(L\mathfrak{g})$ is the quotient of $U_{q}(\widehat{\mathfrak{g}})$ by the ideal generated by the central element $K_{\delta}-1$, where $K_{\delta}$ is defined as in, e.g., [Bec94, §1.3]. Let us recall the “new” Drinfeld presentation of the quantum loop algebra $U_{q}(L\mathfrak{g})$. By [Dri87, Bec94], $U_{q}(L\mathfrak{g})$ is isomorphic to the algebra generated by $x^{\pm}_{i,k},h_{i,l},K^{\pm 1}_{i}$, where $k\in{\mathbb{Z}}$, $l\in{\mathbb{Z}}-\{0\}$ and $i\in\mathbb{I}_{0}$, subject to the following relations:
(2.7) $K_{i}K_{i}^{-1}=$ $\ K_{i}^{-1}K_{i}=1,$
(2.8) $K_{i}K_{j}=$ $\ K_{j}K_{i},$
(2.9) $[h_{i,k},h_{j,l}]=$ $\ 0,$
(2.10) $K_{i}h_{j,k}=$ $\ h_{j,k}K_{i},$
(2.11) $K_{i}x^{\pm}_{j,k}=$ $\ q_{i}^{\pm a_{ji}}x^{\pm}_{j,k}K_{i},$
(2.12) $[h_{i,k},x^{\pm}_{j,l}]=$$\pm\textstyle\frac{[k\cdot a_{ji}]_{q_{i}}}{k}x^{\pm}_{j,k+l},$
(2.13) $x^{\pm}_{i,k+1}x^{\pm}_{j,l}-q_{i}^{\pm a_{ji}}x^{\pm}_{j,l}x^{\pm}_{i,k+1}=$ $\ q_{i}^{\pm a_{ji}}x^{\pm}_{i,k}x^{\pm}_{j,l+1}-x^{\pm}_{j,l+1}x^{\pm}_{i,k},$
(2.14) $[x^{+}_{i,k},x^{-}_{j,l}]=$ $\ \delta_{ij}\textstyle\frac{1}{q_{i}-q_{i}^{-1}}(\phi^{+}_{i,k+l}-\phi^{-}_{i,k+l}),$
[…]
where $\operatorname{Sym}_{k_{1},\cdots,k_{r}}$ denotes symmetrization with respect to the indices $k_{1},\cdots,k_{r}$ and
$\boldsymbol{\phi}^{\pm}_{i}(z)=\sum_{k=0}^{\infty}\phi^{\pm}_{i,\pm k}z^{\pm k}=K_{i}^{\pm 1}\exp\left(\pm(q-q^{-1})\sum_{k=1}^{\infty}h_{i,\pm k}z^{\pm k}\right).$
Let
$\widetilde{U}_{q}(L\mathfrak{g})=U_{q}(L\mathfrak{g})\otimes{\mathbb{C}}[\mathbb{K}_{i}^{\pm 1}\mid i\in\mathbb{I}].$
It can be regarded as a version of the Drinfeld double of $U_{q}(\widehat{\mathfrak{n}}_{+})$. We consider $\widetilde{U}_{q}(L\mathfrak{g})$ as a $Q$-graded algebra with
$\operatorname{deg}^{\mathrm{Dr}}x^{\pm}_{i,k}=\pm\alpha_{i},\quad\operatorname{deg}^{\mathrm{Dr}}h_{i,k}=\operatorname{deg}^{\mathrm{Dr}}\mathbb{K}_{i}=0\qquad(i\in\mathbb{I}_{0}).$
In the following, we will use the abbreviations
$\mathbf{U}=U_{q}(L\mathfrak{g}),\qquad\widetilde{\mathbf{U}}=\widetilde{U}_{q}(L\mathfrak{g}).$
Let $\widetilde{\mathbf{U}}_{+}$ be the subalgebra of $\widetilde{\mathbf{U}}$ spanned by elements of positive degree, i.e., non-negative along all $\alpha_{i}$, and positive along at least one $\alpha_{i}$ ($i\in\mathbb{I}_{0}$).
We will also need to consider other subalgebras of $\widetilde{\mathbf{U}}$ spanned by elements of specified degree. Write $\operatorname{deg}^{\mathrm{Dr}}_{i}$ for degree along the simple root $\alpha_{i}$. Let $Q_{i}$ be the sublattice of $Q$ spanned by all the simple roots except $\alpha_{i}$. Let $\widetilde{\mathbf{U}}_{d_{i}=r,+}$ be the subalgebra of $\widetilde{\mathbf{U}}$ spanned by elements which are both: (i) of degree $r$ along $\alpha_{i}$, and (ii) positive $Q_{i}$-degree, i.e., non-negative along all the simple roots and positive along at least one simple root in $Q_{i}$. That is,
$\widetilde{\mathbf{U}}_{d_{i}=r,+}=\langle y\in\widetilde{\mathbf{U}}\mid\operatorname{deg}^{\mathrm{Dr}}_{i}y=r;\ \forall j\neq i\in\mathbb{I}_{0}\colon\operatorname{deg}^{\mathrm{Dr}}_{j}y\geq 0;\ \exists j\neq i\colon\operatorname{deg}^{\mathrm{Dr}}_{j}y>0\rangle.$
Moreover, set
(2.15) $\widetilde{\mathbf{U}}_{d_{i}\geq r,+}=$ $\ \langle y\in\widetilde{\mathbf{U}}\mid\operatorname{deg}^{\mathrm{Dr}}_{i}y\geq r;\ \forall j\neq i\in\mathbb{I}_{0}\colon\operatorname{deg}^{\mathrm{Dr}}_{j}y\geq 0;\ \exists j\neq i\colon\operatorname{deg}^{\mathrm{Dr}}_{j}y>0\rangle,$
(2.16) $\widetilde{\mathbf{U}}_{\neq i,+}=$ $\ \bigoplus_{r\in{\mathbb{Z}}}\widetilde{\mathbf{U}}_{d_{i}=r,+}=\langle y\in\widetilde{\mathbf{U}}\mid\forall j\neq i\in\mathbb{I}_{0}\colon\operatorname{deg}^{\mathrm{Dr}}_{j}y\geq 0;\ \exists j\neq i\colon\operatorname{deg}^{\mathrm{Dr}}_{j}y>0\rangle,$
(2.17) $\widetilde{\mathbf{U}}_{+}=$ $\ \langle y\in\widetilde{\mathbf{U}}\mid\forall j\in\mathbb{I}_{0}\colon\operatorname{deg}^{\mathrm{Dr}}_{j}y\geq 0,\ \exists j\colon\operatorname{deg}^{\mathrm{Dr}}_{j}y>0\rangle.$
We will also need the following lemma.
Lemma 2.1.
We have
$\operatorname{deg}^{\mathrm{Dr}}e_{0}^{\pm}=\mp\theta.$
Proof.
This follows directly from the explicit description of the images of $E_{0}$ and $F_{0}$ in the Drinfeld realization, see, e.g., [Bec94, Remark 4.7] and [CP94, Theorem 2.2]. ∎
2.2. The braid group action
The Weyl group $W_{0}$ of ${\mathfrak{g}}$ is generated by the simple reflections $s_{i}$ ($i\in\mathbb{I}_{0}$), and acts on $P$ in the usual way. The extended affine Weyl group $\widetilde{W}$ is the semi-direct product $W_{0}\ltimes P$. It contains the affine Weyl group $W=W_{0}\ltimes Q=\langle s_{i}\mid i\in\mathbb{I}\rangle$ as a subgroup, and can also be realized as the semi-direct product
(2.18) $\widetilde{W}\cong\Lambda\ltimes W,$
where $\Lambda=P/Q$ is a finite group of automorphisms of the Dynkin diagram of $\widehat{{\mathfrak{g}}}$. Let $\widetilde{\mathcal{B}}$ be the braid group associated to $\widetilde{W}$.
It is well known [Lus10, §37.1.3] that, for each $i\in\mathbb{I}$, there exists an automorphism $T_{i}$ of $U_{q}(L\mathfrak{g})$ such that $T_{i}(K_{\mu})=K_{s_{i}\mu}$ and
(2.19) $T_{i}(E_{i})=$ $\ -F_{i}K_{i},\quad$ $T_{i}(E_{j})=$ $\ \sum_{r+s=-a_{ji}}(-1)^{r}q_{i}^{-r}E_{i}^{(s)}E_{j}E_{i}^{(r)},$
(2.20) $T_{i}(F_{i})=$ $\ -K_{i}^{-1}E_{i},\quad$ $T_{i}(F_{j})=$ $\ \sum_{r+s=-a_{ji}}(-1)^{r}q_{i}^{r}F_{i}^{(r)}F_{j}F_{i}^{(s)},$
for $\mu\in P$ and $i\neq j$. For each $\lambda\in\Lambda$, there is also an automorphism $T_{\lambda}$ such that $T_{\lambda}(E_{i})=E_{\lambda(i)}$, $T_{\lambda}(F_{i})=F_{\lambda(i)}$ and $T_{\lambda}(K_{i})=K_{\lambda(i)}$. These automorphisms satisfy the relations of the braid group $\widetilde{\mathcal{B}}$. We extend this action to an action on $\widetilde{\mathbf{U}}$ by setting $T_{i}(\mathbb{K}_{\mu})=\mathbb{K}_{s_{i}\mu}$ and $T_{\lambda}(\mathbb{K}_{i})=\mathbb{K}_{\lambda(i)}$. Given a reduced expression $w=\lambda s_{i_{1}}\cdots s_{i_{r}}\in\widetilde{W}$, one defines $T_{w}=T_{\lambda}T_{i_{1}}\cdots T_{i_{r}}$. This is independent of the choice of reduced expression.
2.3. Quantum symmetric pairs of split affine type
We recall the definition of quantum symmetric pair coideal subalgebras of split affine type, i.e., corresponding to Satake diagrams with no black nodes and no involution. More precisely, we consider a “universal” version of these algebras introduced in [LW22, Zha22], also known as universal $\imath$quantum groups.
Definition 2.2.
Let $\widetilde{\mathbf{U}}^{\imath}=\widetilde{\mathbf{U}}^{\imath}(\widehat{{\mathfrak{g}}})$ be the algebra generated by $B_{i}$ and central invertible elements $\mathbb{K}_{i}$ $(i\in\mathbb{I})$ subject to relations
$-q_{i}\mathbb{K}_{i}^{-1}\mathsf{Serre}_{ij}(B_{i},B_{j})=\left\{\begin{array}[]{r l l}0&\mbox{if}&a_{ji}=0,\\ B_{j}&\mbox{if}&a_{ji}=-1,\\ {[2]}_{q_{i}}^{2}[B_{i},B_{j}]&\mbox{if}&a_{ji}=-2,\\ (1+[3]_{q_{i}})(B_{i}^{2}B_{j}+B_{j}B_{i}^{2})\\ -[4]_{q_{i}}(1+[2]_{q_{i}}^{2})B_{i}B_{j}B_{i}+q_{i}^{-1}[3]_{q_{i}}^{2}\mathbb{K}_{i}B_{j}&\mbox{if}&a_{ji}=-3.\end{array}\right.$
It follows from [Kol14, Theorem 8.3] that $Z(\widetilde{\mathbf{U}}^{\imath})={\mathbb{C}}[\mathbb{K}_{i}^{\pm 1}\mid i\in\mathbb{I}]$. Recently, Lu–Wang [LW21] and Zhang [Zha22] found a “Drinfeld-type” presentation of $\widetilde{\mathbf{U}}^{\imath}$.
Definition 2.3.
Let ${}^{\mathrm{Dr}}\widetilde{\mathbf{U}}^{\imath}$ be the $\mathbb{C}$-algebra generated by $H_{i,m}$ and $A_{i,r}$, where $m\geq 1$, $r\in{\mathbb{Z}}$, and invertible central elements $\mathbb{K}_{i}$ ($i\in\mathbb{I}_{0}$), $\mathfrak{C}$, subject to the following relations:
(2.21) $[H_{i,m},H_{j,l}]$ $=0,$
(2.22) $[H_{i,m},A_{j,r}]$ $=\textstyle\frac{[m\cdot a_{ji}]_{q_{i}}}{m}(A_{j,r+m}-A_{j,r-m}\mathfrak{C}^{m}),$
(2.23) $[A_{i,r},A_{j,s}]$ $=0\quad\mbox{if}\quad a_{ji}=0,$
(2.24) $[A_{i,r},A_{j,s+1}]_{q_{i}^{-a_{ji}}}-q_{i}^{-a_{ji}}[A_{i,r+1},A_{j,s}]_{q^{a_{ji}}}$ $=0\quad\mbox{if}\quad i\neq j,$
(2.25) $[A_{i,r},A_{i,s+1}]_{q_{i}^{-2}}-q_{i}^{-2}[A_{i,r+1},A_{i,s}]_{q_{i}^{2}}$ $=q_{i}^{-2}\mathbb{K}_{i}\mathfrak{C}^{r}\Theta_{i,s-r+1}-q_{i}^{-4}\mathbb{K}_{i}\mathfrak{C}^{r+1}\Theta_{i,s-r-1}$
$\quad+q_{i}^{-2}\mathbb{K}_{i}\mathfrak{C}^{s}\Theta_{i,r-s+1}-q_{i}^{-4}\mathbb{K}_{i}\mathfrak{C}^{s+1}\Theta_{i,r-s-1},$
and Serre relations (see [Zha22, (3.27)-(3.28)]), where $m,n\geq 1$; $r,s,r_{1},r_{2}\in{\mathbb{Z}}$ and
$1+\sum_{m=1}^{\infty}(q-q^{-1})\Theta_{i,m}z^{m}=\exp\left((q-q^{-1})\sum_{m=1}^{\infty}H_{i,m}z^{m}\right).$
By convention, $\Theta_{i,0}=(q-q^{-1})^{-1}$ and $\Theta_{i,m}=0$ for $m\leq-1$.
By [Zha22, Theorem 3.4], there is an algebra isomorphism
[…]
This isomorphism and its inverse can be described in terms of the braid group action recalled below in §2.4. For details, see [Zha22, (3.1)-(3.3)]. In particular, it sends
(2.27) $B_{i}\mapsto A_{i,0},\quad\mathbb{K}_{i}\mapsto\mathbb{K}_{i},\quad\mathbb{K}_{\delta}\mapsto\mathfrak{C},$
for $i\in\mathbb{I}_{0}$.
2.4. The QSP braid group action
By [Zha22, Lemma 2.7], for each $i\in\mathbb{I}$, there exists an automorphism $\mathbf{T}_{i}$ of $\widetilde{\mathbf{U}}^{\imath}$ such that $\mathbf{T}_{i}(\mathbb{K}_{\mu})=\mathbb{K}_{s_{i}\mu}$ and
$\mathbf{T}_{i}(B_{j})=\left\{\begin{array}[]{l l}\mathbb{K}_{j}^{-1}B_{j}&\mbox{if }i=j,\\[2.0pt] B_{j}&\mbox{if }a_{ji}=0,\\[2.0pt] B_{j}B_{i}-q_{i}B_{i}B_{j}&\mbox{if }a_{ji}=-1,\\[2.0pt] {[2]_{q_{i}}^{-1}}(B_{j}B_{i}^{2}-q_{i}[2]_{q_{i}}B_{i}B_{j}B_{i}+q_{i}^{2}B_{i}^{2}B_{j})+B_{j}\mathbb{K}_{i}&\mbox{if }a_{ji}=-2,\\[2.0pt] {[3]}_{q_{i}}^{-1}[2]_{q_{i}}^{-1}\big{(}B_{j}B_{i}^{3}-q_{i}[3]_{q_{i}}B_{i}B_{j}B_{i}^{2}+q^{2}[3]_{q_{i}}B_{i}^{2}B_{j}B_{i}\\[2.0pt] -q_{i}^{3}B_{i}^{3}B_{j}+q_{i}^{-1}[B_{j},B_{i}]_{q_{i}^{3}}\mathbb{K}_{i}\big{)}+[B_{j},B_{i}]_{q_{i}}\mathbb{K}_{i}&\mbox{if }a_{ji}=-3,\end{array}\right.$
for $\mu\in{\mathbb{Z}}\mathbb{I}$ and $j\in\mathbb{I}$. For each $\lambda\in\Lambda$, there is also an automorphism $\mathbf{T}_{\lambda}$ such that $\mathbf{T}_{\lambda}(B_{i})=B_{\lambda(i)}$ and $\mathbf{T}_{\lambda}(\mathbb{K}_{i})=\mathbb{K}_{\lambda(i)}$. These automorphisms define an action of the braid group $\widetilde{\mathcal{B}}$. We will refer to this action as the QSP braid group action to distinguish it from Lusztig’s braid group action from §2.2.
We will need the following lemma.
Lemma 2.4.
We have $T_{w}(B_{i})=B_{wi}$, $\mathbf{T}_{w}(B_{i})=B_{wi}$, for $i\in\mathbb{I}$ and $w\in\widetilde{W}$ such that $wi\in\mathbb{I}$.
Proof.
See [Lus10] and [Zha22, Lemma 2.10]. ∎
2.5. Coideal structures
By [Kol14, Theorem 7.1], for any $\mathbf{s}=(s_{0},\cdots,s_{n})\in{\mathbb{C}}^{n+1}$, there exists an injective algebra homomorphism
(2.28) $\eta_{\mathbf{s}}\colon\widetilde{\mathbf{U}}^{\imath}\hookrightarrow\widetilde{\mathbf{U}},\quad B_{i}\mapsto e_{i}^{-}-q_{i}^{-2}\mathbb{K}_{i}e_{i}^{+}K_{i}^{-1}+s_{i}K_{i}^{-1},\quad\mathbb{K}_{i}\mapsto\mathbb{K}_{i}\quad(i\in\mathbb{I}).$
Given $\mathbf{c}=(c_{0},\cdots,c_{n})\in{\mathbb{C}}^{n+1}$, let $\mathbf{U}^{\imath}_{\mathbf{c}}$ be the quotient of $\widetilde{\mathbf{U}}^{\imath}$ by the two-sided ideal generated by $q_{i}^{-2}\mathbb{K}_{i}-c_{i}$. The induced map $\eta_{\mathbf{c},\mathbf{s}}\colon\mathbf{U}^{\imath}_{\mathbf{c}}\hookrightarrow\mathbf{U}$ gives $\mathbf{U}^{\imath}_{\mathbf{c}}$ the structure of a right coideal subalgebra of $\mathbf{U}$, with coproduct $\Delta_{\mathbf{c},\mathbf{s}}=\Delta\circ\eta_{\mathbf{c},\mathbf{s}}$. If $\mathbf{s}=(0,\cdots,0)$, we abbreviate $\eta=\eta_{\mathbf{s}}$ and $\eta_{\mathbf{c}}=\eta_{\mathbf{c},\mathbf{s}}$. Explicitly,
(2.29) $\Delta_{\mathbf{c},\mathbf{s}}(B_{i})=1\otimes\eta_{\mathbf{c}}(B_{i})+\eta_{\mathbf{c},\mathbf{s}}(B_{i})\otimes K_{i}^{-1}.$
2.6. Rank one subalgebras
Let us recall how the braid group actions can be used to obtain rank one subalgebras in $\widetilde{\mathbf{U}}$ and $\widetilde{\mathbf{U}}^{\imath}$. For $i\in\mathbb{I}_{0}$, let $\omega^{\prime}_{i}=\omega_{i}s_{i}$, and let $\mathbf{U}_{[i]}$ be the subalgebra of $\mathbf{U}$ generated by
$E_{i},\ F_{i},\ K_{i}^{\pm 1},\ T_{\omega^{\prime}_{i}}(E_{i}),\ T_{\omega^{\prime}_{i}}(F_{i}),\ T_{\omega^{\prime}_{i}}(K_{i}^{\pm 1}).$
By [Bec94, Proposition 3.8], there is an algebra isomorphism $\iota_{i}\colon U_{q}(L\mathfrak{sl}_{2})\to\mathbf{U}_{[i]}$ sending $q\mapsto q_{i}$ and
$E_{1}\mapsto E_{i},\ F_{1}\mapsto F_{i},\ K_{1}^{\pm 1}\mapsto K_{i}^{\pm 1},\quad E_{0}\mapsto T_{\omega^{\prime}_{i}}(E_{i}),\ F_{0}\mapsto T_{\omega^{\prime}_{i}}(F_{i}),\ K_{0}^{\pm 1}\mapsto T_{\omega^{\prime}_{i}}(K_{i}^{\pm 1}).$
Let $\widetilde{\mathbf{U}}_{[i]}$ be the subalgebra of $\widetilde{\mathbf{U}}$ generated by $\mathbf{U}_{[i]}$, $\mathbb{K}_{i}^{\pm 1}$ and $(\mathbb{K}_{\delta}\mathbb{K}_{i}^{-1})^{\pm 1}$. The algebra isomorphism $\iota_{i}$ extends to an isomorphism $\iota_{i}\colon\widetilde{U}_{q}(L\mathfrak{sl}_{2})\to\widetilde{\mathbf{U}}_{[i]}$ sending $\mathbb{K}_{1}\mapsto K_{i}$ and $\mathbb{K}_{0}\mapsto\mathbb{K}_{\delta}\mathbb{K}_{i}^{-1}$.
Moreover, for $i\in\mathbb{I}_{0}$, let $\widetilde{\mathbf{U}}^{\imath}_{[i]}$ be the subalgebra of $\widetilde{\mathbf{U}}^{\imath}$ generated by $B_{i}$, $\mathbf{T}_{\omega^{\prime}_{i}}(B_{i})$, $\mathbb{K}_{i}^{\pm 1}$ and $(\mathbb{K}_{\delta}\mathbb{K}_{i}^{-1})^{\pm 1}$. By [LW21, Proposition 3.9], there is an algebra isomorphism $\iota\colon\widetilde{\mathbf{U}}^{\imath}(\widehat{\mathfrak{sl}}_{2})\to\widetilde{\mathbf{U}}^{\imath}_{[i]}$ sending $q\mapsto q_{i}$ and
(2.30) $B_{1}\mapsto B_{i},\quad B_{0}\mapsto\mathbf{T}_{\omega^{\prime}_{i}}(B_{i}),\quad\mathbb{K}_{1}\mapsto\mathbb{K}_{i},\quad\mathbb{K}_{0}\mapsto\mathbb{K}_{\delta}\mathbb{K}_{i}^{-1}.$
We also note that $A_{-1}$ is sent to $\mathbf{T}_{\omega_{i}}(B_{i})$.
2.7. Factorization and coproduct in rank one
Let $\mathcal{H}$ denote the commutative subalgebra of $\widetilde{\mathbf{U}}$ generated by the coefficients of the series $\boldsymbol{\phi}^{\pm}_{i}(z)$. We will refer to these coefficients as the Drinfeld–Cartan operators, and to $\mathcal{H}$ as the Drinfeld–Cartan subalgebra of $\widetilde{\mathbf{U}}$. Similarly, we refer to the coefficients of the series
$\boldsymbol{\grave{\Theta}}_{i}(z)=(q_{i}-q_{i}^{-1})\frac{1-q_{i}^{-2}\mathfrak{C}z^{2}}{1-\mathfrak{C}z^{2}}\sum_{r\geq 0}\Theta_{i,r}z^{r}$
as $\imath$Drinfeld–Cartan operators, and to the commutative subalgebra $\mathcal{H}^{\imath}$ generated by them as the $\imath$Drinfeld–Cartan subalgebra of $\widetilde{\mathbf{U}}^{\imath}$. It is natural to ask how the two Drinfeld–Cartan subalgebras are related under the monomorphism (2.28). This question was answered in [Prz23] for quantum symmetric pairs of split affine type $\mathsf{A}$. Here we will need the rank one case of that result, i.e., when $\widetilde{\mathbf{U}}^{\imath}$ is isomorphic to the (universal) $q$-Onsager algebra.
Theorem 2.5 ([Prz23, Theorem 7.5]).
Let $\widehat{{\mathfrak{g}}}=\widehat{\mathfrak{sl}}_{2}$. The series $\boldsymbol{\grave{\Theta}}(z)$ admits the following factorization
(2.31) $\eta_{\mathbf{s}}(\boldsymbol{\grave{\Theta}}(z))\equiv$ $\ \xi_{\mathbf{s}}(\boldsymbol{\grave{\Theta}}(z))\cdot\boldsymbol{\phi}^{-}(z^{-1})\boldsymbol{\phi}^{+}(\mathfrak{C}z)\quad\mod{\widetilde{\mathbf{U}}_{+}}[\negthinspace[z]\negthinspace],$
(2.32) $\Delta_{\mathbf{s}}(\boldsymbol{\grave{\Theta}}(z))\equiv$ $\ \eta_{\mathbf{s}}(\boldsymbol{\grave{\Theta}}(z))\otimes\eta(\boldsymbol{\grave{\Theta}}(z))\quad\mod{\widetilde{\mathbf{U}}\otimes\widetilde{\mathbf{U}}_{+}}[\negthinspace[z]\negthinspace],$
where $\xi_{\mathbf{s}}=\varepsilon\circ\eta_{\mathbf{s}}$.
3. Reduced expressions for fundamental weights
The goal of this section is to obtain explicit reduced expressions for the extended affine Weyl group elements corresponding to the fundamental weights in types $\mathsf{A}_{n}^{(1)},\mathsf{B}_{n}^{(1)}$, $\mathsf{C}_{n}^{(1)},\mathsf{D}_{n}^{(1)}$.
First, we need to introduce some notation. Let $w_{0}$ be the longest element of the finite Weyl group $W_{0}$ of ${\mathfrak{g}}$. Given $i\in\mathbb{I}_{0}$, let $w_{i}$ be the longest element of the Weyl group corresponding to the Dynkin diagram of ${\mathfrak{g}}$ with the node labelled by $i$ removed. Given $1\leq k<l$, let
$[k,l]=s_{k}s_{k+1}\cdots s_{l},\quad[l,k]=[k,l]^{-1}.$
Given the highest root $\theta=\sum_{i\in\mathbb{I}_{0}}c_{i}\alpha_{i}$, let $J=\{i\in\mathbb{I}_{0}\mid c_{i}=1\}$.
We will use the following two results.
Proposition 3.1.
There is a bijection
$J\ \to\ \Lambda{-}\{1\},\qquad i\mapsto\omega_{i}w_{i}w_{0}.$
Proof.
See [Bou02, Ch. VI §2.3, Proposition 6] or [IM65, Proposition 1.18]. ∎
Proposition 3.2.
The length of the fundamental weights is given by
$\ell(\omega_{i})=\sum_{\beta\in\Delta^{+}}(\beta,\omega_{i}),$
where $\Delta^{+}$ are the positive roots of ${\mathfrak{g}}$.
Proof.
See [IM65, Proposition 1.23]. ∎
$\mathsf{A}_{n}^{(1)}$
[…]
$\mathsf{B}_{2}^{(1)}$
2222111100
$\mathsf{B}_{n}^{(1)}$, $n\geq 3$
[…]
$\mathsf{C}_{n}^{(1)}$, $n\geq 3$
[…]
$\mathsf{D}_{n}^{(1)}$, $n\geq 4$
[…]
Table 2. Affine Dynkin diagrams of types $\mathsf{A}_{n}^{(1)}$, $\mathsf{B}_{n}^{(1)}$, $\mathsf{C}_{n}^{(1)}$, $\mathsf{D}_{n}^{(1)}$.
3.1. Type $\mathsf{A}_{n}^{(1)}$
The highest root is $\theta=\alpha_{1}+\alpha_{2}+\cdots+\alpha_{n}$ and $J=\mathbb{I}_{0}$. The fundamental group $\Lambda$ is cyclic of order $n+1$. It is generated by, e.g., the automorphism $\pi$ of the affine Dynkin diagram sending $i\mapsto i+1\mod n+1$.
Proposition 3.3.
Let $\widetilde{W}$ be the extended affine Weyl group of type $A_{n}^{(1)}$, for $n\geq 1$. Then
$\omega_{i}=\pi^{i}[n-i+1,n]\cdots[2,i+1][1,i].$
This is a reduced expression.
Proof.
See [Prz23, Proposition 9.2]. ∎
3.2. Type $\mathsf{D}_{n}^{(1)}$
The highest root is $\theta=\alpha_{1}+\alpha_{n-1}+\alpha_{n}+2(\alpha_{2}+\cdots+\alpha_{n-2})$ and $J=\{1,n{-}1,n\}$. The fundamental group $\Lambda=\{1,\pi_{1},\pi_{n-1},\pi_{n}\}$ depends on the parity of $n$. If $n$ is even then $\Lambda\cong{\mathbb{Z}}/2{\mathbb{Z}}\times{\mathbb{Z}}/2{\mathbb{Z}}$, and if $n$ is odd then $\Lambda\cong{\mathbb{Z}}/4{\mathbb{Z}}$. Explicit descriptions of the corresponding diagram automorphisms can be found in [Bou02, Ch. VI §4.8]. Given $21\leq m\leq n-$ 2, set
$r_{m}=s_{n}[m,n{\shortminus}1][m{\shortminus}1,n{\shortminus}2].$
Proposition 3.4.
Let $\widetilde{W}$ be the extended affine Weyl group of type $\mathsf{D}_{n}^{(1)}$, for $n\geq 4$. If $i$ is even, then
$\omega_{i}=\left\{\begin{array}[]{r l }\big{(}s_{0}[2,n{\shortminus}1][1,n{\shortminus}2]s_{n}\big{)}^{\frac{i}{2}}[{n{\shortminus}i},{n{\shortminus}1}]\cdots[1,i]&\quad\mbox{if }\ 2\leq i\leq n{\shortminus}2,\\ \pi_{n{\shortminus}1}r_{n{\shortminus}2}r_{n{\shortminus}4}\cdots r_{3}s_{n}[1,{n{\shortminus}1}]&\quad\mbox{if }\ i=n{\shortminus}1,\\ \pi_{n}r_{n{\shortminus}2}r_{n{\shortminus}4}\cdots r_{2}s_{n}&\quad\mbox{if }\ i=n.\\ \end{array}\right.$
If $i$ is odd, then
$\omega_{i}=\left\{\begin{array}[]{r l }\pi_{1}[1,n{\shortminus}2]s_{n}\big{(}s_{0}[2,n{\shortminus}1][1,n{\shortminus}2]s_{n}\big{)}^{\frac{i{\shortminus}1}{2}}[{n{\shortminus}i},{n{\shortminus}1}]\cdots[1,i]&\quad\mbox{if }\ 1\leq i\leq n{\shortminus}2,\\ \pi_{n{\shortminus}1}s_{n{\shortminus}1}s_{n{\shortminus}2}r_{n{\shortminus}3}r_{n{\shortminus}5}\cdots r_{3}s_{n}[1,{n{\shortminus}1}]&\quad\mbox{if }\ i=n{\shortminus}1,\\ \pi_{n}s_{n{\shortminus}1}s_{n{\shortminus}2}r_{n{\shortminus}3}r_{n{\shortminus}5}\cdots r_{2}s_{n}&\quad\mbox{if }\ i=n.\\ \end{array}\right.$
Moreover, the formulae above yield reduced expressions.
Proof.
If $i\in J=\{1,n{-}1,n\}$, one may use Proposition 3.1 and argue in a similar way as in Proposition 3.3. For more general $i\in\mathbb{I}_{0}$, consider $\widetilde{W}$ as the group of affine transformations of ${\mathbb{R}}^{n}$. We show that elements on the RHS of the equations above act in the same way as $\omega_{i}$. Indeed, the element $[{n{\shortminus}i},{n{\shortminus}1}]\cdots[1,i]$ acts as a cyclic permutation of the coordinates (shifting the indices by $n-i$). It is also easy to explicitly verify that $\big{(}s_{0}[2,n{\shortminus}1][1,n{\shortminus}2]s_{n}\big{)}^{\frac{i}{2}}$ (in the even case) and $\pi_{1}\big{(}[1,n{\shortminus}2]s_{n}s_{0}[2,n{\shortminus}1]\big{)}^{\frac{i{\shortminus}1}{2}}[1,n{\shortminus}2]s_{n}$ (in the odd case) act as cyclic permutations of the coordinates (shifting the indices by $i$) followed by an affine shift by $\omega_{i}$.
We will prove the reducedness of our expressions for $\omega_{i}$ by induction on $n$. The base case of $\mathsf{D}_{4}$ can be checked directly. Now assume that $n\geq 5$, and that the result is true for $\mathsf{D}_{n-1}$. The set of positive roots of type $\mathsf{D}_{n}$ ($n\geq 5$) is the union of the set of positive roots of type $\mathsf{D}_{n-1}$ (with the indices shifted by $1$), and the set consisting of the roots
(3.1) $\sum_{j=1}^{l}\alpha_{j}\ (1\leq l\leq n),\quad\alpha_{n}+\sum_{j=1}^{n-2}\alpha_{j},\quad\sum_{j=1}^{n}\alpha_{j}+\sum_{j=2}^{l}\alpha_{n-j}\ (1\leq l\leq n-2).$
Let $21\leq i\leq n-$ 2. We will show that $\ell(\omega_{i})=i(2n-i-1)$. By Proposition 3.2, it suffices to check that the sum of the coefficients of $\alpha_{i}$ in positive roots of type $\mathsf{D}_{n}$ (written in terms of the basis of simple roots) is equal to $i(2n-i-1)$. This sum is equal to the sum of the coefficients of $\alpha_{i-1}$ in the positive roots of type $\mathsf{D}_{n-1}$ and the sum of the coefficients of $\alpha_{i}$ in the roots in (3.1). By induction, the former is equal to $(i-1)(2n-2-i)$. Therefore, the total sum is $(i-1)(2n-2-i)+(n-i+1)+1+(n-3)+(i-1)=i(2n-i-1)$.
Next, let $i=n$. We will check that $\ell(\omega_{n})=n(n-1)/2$. By induction, the sum of the coefficients of $\alpha_{n-1}$ in the positive roots of type $\mathsf{D}_{n-1}$ is $(n-1)(n-2)/2$. Therefore the total sum of the coefficients of $\alpha_{n}$ in the positive roots of type $\mathsf{D}_{n}$ is $(n-1)(n-2)/2+n-1=n(n-1)/2$.
Finally, let $i=n-1$. We will check that $\ell(\omega_{n})=n(n-1)/2$. By induction, the sum of the coefficients of $\alpha_{n-2}$ in the positive roots of type $\mathsf{D}_{n-1}$ is $(n-1)(n-2)/2$. Therefore the total sum of the coefficients of $\alpha_{n-1}$ in the positive roots of type $\mathsf{D}_{n}$ is $(n-1)(n-2)/2+n-1=n(n-1)/2$.
In each of the three cases, the length of $\omega_{i}$ equals the number of letters in the expressions for $\omega_{i}$ we found, implying that these expressions are reduced. ∎
3.3. Type $\mathsf{B}_{n}^{(1)}$
The highest root is $\theta=\alpha_{1}+2\sum_{i=2}^{n}\alpha_{i}$ and $J=\{1\}$. The fundamental group is $\Lambda=\{1,\pi_{1}\}$, where $\pi_{1}$ interchanges $\alpha_{0}$ and $\alpha_{1}$, and fixes the other $\alpha_{j}$’s.
Proposition 3.5.
Let $\widetilde{W}$ be the extended affine Weyl group of type $\mathsf{B}_{n}^{(1)}$, for $n\geq 2$. If $i$ is odd, then [^1] If $i=n$, the expression below reads $\pi_{1}[1,n]\big{(}s_{0}[2,n][1,n]\big{)}^{\frac{n-1}{2}}$.
$\omega_{i}=\pi_{1}[1,n]\big{(}s_{0}[2,n][1,n]\big{)}^{\frac{i-1}{2}}[{n{\shortminus}i},{n{\shortminus}1}]\cdots[1,i].$
If $i$ is even, then [^2] If $i=n$, the expression below reads $\big{(}s_{0}[2,n][1,n]\big{)}^{\frac{n}{2}}$.
$\omega_{i}=\big{(}s_{0}[2,n][1,n]\big{)}^{\frac{i}{2}}[{n{\shortminus}i},{n{\shortminus}1}]\cdots[1,i].$
Moreover, the formulae above yield reduced expressions.
Proof.
Given $I\subset\{1,n\}$, let $w_{0}^{I}$ denote the longest word in the parabolic subgroup $W_{I}$.
First consider the case of $i=1$. By Proposition 3.1, $\omega_{1}=\pi_{1}w_{0}w_{1}$. We have $w_{0}=w_{0}^{[1,n-1]}(s_{n}(s_{n-1}s_{n})\cdots[2,n])^{-1}$ and $w_{1}=s_{n}(s_{n-1}s_{n})\cdots[2,n]w_{0}^{[2,n-1]}$. Therefore $\pi_{1}w_{0}w_{1}=\pi_{1}w_{0}^{[1,n-1]}w_{0}^{[2,n-1]}=\pi_{1}[1,n][n-1,1]$.
For $i\neq 1$, we use the same strategy as in the proof of Proposition 3.4, i.e., show that the expressions we found define the same affine transformations of ${\mathbb{R}}^{n}$. More precisely, we need to show that each $\omega_{i}$ (as an element of the extended affine Weyl group of type $\mathsf{B}_{n}$) acts as a translation by the $i$-th fundamental weight of type $\mathsf{C}_{n}$. Let us first recall some facts about the root systems of types $\mathsf{B}_{n}$ and $\mathsf{C}_{n}$. The simple roots in the root system of type $\mathsf{B}_{n}$ are
$\alpha_{j}=e_{j}-e_{j+1}\ (1\leq j\leq n-1),\ \ \alpha_{n}=e_{n}.$
We have $\theta=\alpha_{1}+2\sum_{i=2}^{n}\alpha_{i}=e_{1}+e_{2}=\theta^{\vee}$,
(3.2) $s_{j}(e_{j})=$ $\ e_{j+1}$ $(1\leq j\leq n-1),\quad$ $s_{n}(e_{n})$ $=\ -e_{n},$
(3.3) $s_{j}(e_{j+1})=$ $\ e_{j}$ $(1\leq j\leq n-1),\quad$ $s_{0}(e_{j})$ $=\ t_{\theta^{\vee}}r_{\theta}(e_{j}),$
where $t_{\theta^{\vee}}$ is the translation by ${\theta^{\vee}}$; $r_{\theta}$ is the reflection by $\theta$; and in all other cases $s_{j}(e_{j^{\prime}})=e_{j^{\prime}}$. Finally, recall that the fundamental weights of type $\mathsf{C}_{n}$ are $\omega_{i}^{*}=\sum_{j=1}^{i}e_{j}$ ($1\leq i\leq n$).
Observe that $[n-i,n-1]\cdots[2,i+1][1,i]$ acts by cyclically shifting the coordinates by $n-i$. Let $i$ be even. We have
$(s_{0}[2,n][1,n])^{\frac{i}{2}}[n-i,n-1]\cdots[2,i+1][1,i](e_{j})=(s_{0}[2,n][1,n])^{\frac{i}{2}}(e_{n-i+j})$
$=(s_{0}[2,n][1,n])^{\frac{i}{2}-1}t_{\theta}(e_{j-i+2})=(s_{0}[2,n][1,n])^{\frac{i}{2}-1}(e_{j-i+2}+e_{1}+e_{2})$
$=(s_{0}[2,n][1,n])^{\frac{i}{2}-2}(e_{j-i+4}+e_{1}+e_{2}+e_{3}+e_{4})=e_{j}+\sum_{r=1}^{i}e_{r}=e_{j}+\omega_{i}^{*}.$
Let $i$ be odd. We have
$\pi_{1}([1,n]s_{0}[2,n])^{\frac{i-1}{2}}[1,n][n-i,n-1]\cdots[2,i+1][1,i](e_{j})$
$=\pi_{1}([1,n]s_{0}[2,n])^{\frac{i-1}{2}}[1,n](e_{n-i+j})=\pi_{1}[1,n](s_{0}[2,n][1,n])^{\frac{i-1}{2}}(e_{n-i+j})$
$=\pi_{1}[1,n](s_{0}[2,n][1,n])^{\frac{i-3}{2}}(e_{j-i+2}+e_{1}+e_{2})=\pi_{1}[1,n](e_{j-1}+\sum_{r=1}^{i-1}e_{r})$
$=t_{\omega_{1}^{*}}[1,n-1](e_{j-1}+\sum_{r=1}^{i-1}e_{r})=t_{\omega_{1}^{*}}(e_{j}+\sum_{r=2}^{i}e_{r})=e_{j}+\omega_{i}^{*},$
where in the passage from the third to the fourth lines we used the already established formula for $\omega_{1}$.
We will now prove, by induction on $n$, that the expressions for $\omega_{i}$ we produced are reduced. The base case $n=2$ can be checked directly. So assume that $n\geq 3$ and the result is true for $\mathsf{B}_{n-1}$. The set of positive roots of type $\mathsf{B}_{n}$ ($n\geq 3$) is the union of the set of positive roots of type $\mathsf{B}_{n-1}$ (with indices shifted by $1$) and the set consisting of the roots
(3.4) $\sum_{j=1}^{l}\alpha_{j}\ (1\leq l\leq n),\quad\sum_{j=1}^{n}\alpha_{j}+\sum_{j=1}^{l}\alpha_{n-j+1}\ (1\leq l\leq n-1).$
Let $1\leq i\leq n$. We will check that $\ell(\omega_{i})=i(2n-i)$. By Proposition 3.2, it suffices to check that the sum of the coefficients of $\alpha_{i}$ in positive roots of type $\mathsf{B}_{n}$ is equal to $i(2n-i)$. This sum is equal to the sum of the coefficients of $\alpha_{i-1}$ in positive roots of type $\mathsf{B}_{n-1}$ and the sum of the coefficients of $\alpha_{i}$ in the positive roots in (3.4). By induction, the former is equal to $(i-1)(2n-2-i+1)$. Therefore, the total sum is $(i-1)(2n-2-i+1)+(n-i+1)+(n-1)+(i-1)=i(2n-i)$. This is also the length of the expressions for $\omega_{i}$ we found, implying they are indeed reduced. ∎
3.4. Type $\mathsf{C}_{n}^{(1)}$
The highest root is $\theta=\alpha_{n}+2\sum_{i=1}^{n-1}\alpha_{i}$ and $J=\{n\}$. The fundamental group is $\Lambda=\{1,\pi_{n}\}$, where $\pi_{n}$ interchanges $\alpha_{k}$ and $\alpha_{n-k}$.
Proposition 3.6.
Let $\widetilde{W}$ be the extended affine Weyl group of type $\mathsf{C}_{n}^{(1)}$, for $n\geq 2$. If $i=n$, then
(3.5) $\omega_{n}=\pi_{n}s_{n}[n-1,n]\cdots[1,n].$
If $i<n$, then
$\omega_{i}=(s_{0}[1,n])^{i}[n-i,n-1]\cdots[2,i+1][1,i].$
Moreover, the formulae above yield reduced expressions.
Proof.
We follow the same strategy as in the proofs for types $\mathsf{B}_{n}^{(1)}$ and $\mathsf{D}_{n}^{(1)}$. The simple roots in the root system of type $\mathsf{C}_{n}$ are
$\alpha_{j}=e_{j}-e_{j+1}\ (1\leq j\leq n-1),\ \ \alpha_{n}=2e_{n}.$
The highest root is $\theta=2e_{1}=e_{n}+2\sum_{i=1}^{n-1}e_{i}$ and $\theta^{\vee}=e_{1}$. We have
(3.6) $s_{j}(e_{j})=$ $\ e_{j+1}$ $(1\leq j\leq n-1),\quad$ $s_{n}(e_{n})=$ $\ -e_{n},$
(3.7) $s_{j}(e_{j+1})=$ $\ e_{j}$ $(1\leq j\leq n-1),\quad$ $s_{0}(e_{j})=$ $\ t_{\theta^{\vee}}r_{\theta}(e_{j}),$
and in all other cases $s_{j}(e_{j^{\prime}})=e_{j^{\prime}}$. Moreover, the fundamental weights of the dual root system $\mathsf{B}_{n}$ are $\omega_{i}^{*}=\sum_{j=1}^{i}e_{j}$ ($i<n$), and $\omega_{n}^{*}=\frac{1}{2}\sum_{j=1}^{n}e_{j}$.
The case of $i=n$ can simply be handled using Proposition 3.1, i.e., $\omega_{n}=\pi_{n}w_{0}w_{n}=\pi_{n}s_{n}(s_{n-1}s_{n})\cdots[1,n]$. Next, suppose that $i<n$. We have
$(s_{0}[1,n])^{i}[n-i,n-1]\cdots[2,i+1][1,i](e_{j})=(s_{0}[1,n])^{i}(e_{n-i+j}).$
If $j=i$, then
$(s_{0}[1,n])^{i}(e_{n})=(s_{0}[1,n])^{i-1}t_{\theta^{\vee}}r_{\theta}(-e_{1})=(s_{0}[1,n])^{i-1}(2e_{1})$
$=(s_{0}[1,n])^{i-2}(e_{1}+2e_{2})=(s_{0}[1,n])^{i-3}(e_{1}+e_{2}+2e_{3})=e_{i}+\sum_{r=1}^{i}e_{r}=\omega_{i}^{*}+e_{i}.$
If $i<j$, then $j-i+1\geq 2$ and
$(s_{0}[1,n])^{i}(e_{n-i+j})=(s_{0}[1,n])^{i}(e_{j-i})=(s_{0}[1,n])^{i-1}t_{\theta^{\vee}}r_{\theta}(e_{j-i+1})$
$=(s_{0}[1,n])^{i-1}(e_{1}+e_{j-i+1})=(s_{0}[1,n])^{i-2}(e_{1}+e_{2}+e_{j-i+2})=e_{j}+\sum_{r=1}^{i}=e_{j}+\omega_{i}^{*}.$
If $j<i$, then
$(s_{0}[1,n])^{i}(e_{n-i+j})=(s_{0}[1,n])^{i-1}(e_{1}+e_{n-i+j+1})=(s_{0}[1,n])^{i-2}(e_{1}+e_{2}+e_{n-i+j+2})$
$=(s_{0}[1,n])^{j}(e_{n}+\sum_{r=1}^{i-j}e_{r})=(s_{0}[1,n])^{j-1}(e_{1}+\sum_{r=1}^{i-j+1}e_{r})=e_{j}+\sum_{r=1}^{i}e_{r}=e_{j}+\omega_{i}^{*}.$
We will now prove, by induction on $n$, that the expressions for $\omega_{i}$ we found are reduced. The base case of $\mathsf{C}_{2}$ can be checked directly. So assume that $n\geq 3$ and the result is true for $\mathsf{C}_{n-1}$. The set of positive roots of type $\mathsf{C}_{n}$ ($n\geq 3$) is the union of the set of positive roots of type $\mathsf{C}_{n-1}$ (with the indices by $1$) and the set consisting of the roots
(3.8) $\sum_{j=1}^{l}\alpha_{j}\ (1\leq l\leq n),\quad\sum_{j=1}^{n}\alpha_{j}+\sum_{j=1}^{l}\alpha_{n-j}\ (1\leq l\leq n-1).$
We will check that $\ell(\omega_{i})=i(n+1)$ if $i<n$ and $\ell(\omega_{n})=n(n+1)/2$. By Proposition 3.2, it suffices to check that the sum of coefficients of $\alpha_{i}$ in positive roots of type $\mathsf{C}_{n}$ is equal to $i(n-1)$ if $i<n$, and to $n(n+1)/2$ if $i=n$. This sum is equal to the sum of the coefficients of $\alpha_{i-1}$ in positive roots of type $\mathsf{C}_{n-1}$ and the sum of the coefficients of $\alpha_{i}$ in the positive roots in (3.8).
Let $i=n$. By induction hypothesis, the sum of the coefficients of $\alpha_{n-1}$ in positive roots of type $\mathsf{C}_{n-1}$ is equal to $(n-1)n/2$. Therefore, the total sum is $(n-1)n/2+n=n(n+1)/2$. Now let $i<n$. By induction hypothesis, the sum of the coefficients of $\alpha_{i-1}$ in positive roots of type $\mathsf{C}_{n-1}$ is equal to $(i-1)n$. Therefore, the total sum is $(i-1)n+(n-i+1)+(n-1)+i=i(n+1)$, as claimed. This is also the length of the expressions for $\omega_{i}$ we found, implying they are indeed reduced. ∎
4. Auxiliary calculations - root combinatorics
We will need the following lemmas to compute the braid group action in §7, with the aid of Lemma 2.4. Their proofs are based on straightforward calculations, so we omit most of the details.
4.1. Type $\mathsf{D}_{n}^{(1)}$
We consider the cases of weights $\omega_{1},\cdots,\omega_{n-2}$ and $\omega_{n-1},\omega_{n}$ separately.
4.1.1. Weights $\omega_{1},\cdots,\omega_{n-2}$
Lemma 4.1.
We have
(4.6) $s_{0}[2,n{\shortminus}1][1,n{\shortminus}2]s_{n}\cdot\alpha_{j}=$$\ \left\{\begin{array}[]{r l }\alpha_{0}+\alpha_{1}+2\alpha_{2}+\alpha_{3}&\quad\mbox{if }\ j=0,\\ \alpha_{j+2}&\quad\mbox{if }\ 1\leq j\leq n-3,\\ \alpha_{1}&\quad\mbox{if }\ j=n-1,\\ \alpha_{0}+\alpha_{2}+\cdots+\alpha_{n-2}+\alpha_{n}&\quad\mbox{if }\ j=n-2,\\ -\theta-2\alpha_{0}&\quad\mbox{if }\ j=n,\end{array}\right.$
(4.12) $\pi_{1}[1,n{\shortminus}2]s_{n}\cdot\alpha_{j}=$$\ \left\{\begin{array}[]{r l }\alpha_{0}+\alpha_{1}+\alpha_{2}&\quad\mbox{if }\ j=0,\\ \alpha_{j+1}&\quad\mbox{if }\ 1\leq j\leq n-3,\\ \alpha_{0}+\alpha_{2}+\cdots+\alpha_{n-2}+\alpha_{n-1}&\quad\mbox{if }\ j=n-1,\\ \alpha_{n}&\quad\mbox{if }\ j=n-2,\\ -(\alpha_{0}+\alpha_{2}+\cdots+\alpha_{n-2}+\alpha_{n})&\quad\mbox{if }\ j=n,\end{array}\right.$
(4.13) $(s_{0}[2,n{\shortminus}1][1,n{\shortminus}2]s_{n})^{2}\cdot\alpha_{n-2}=$ $\ \alpha_{2},$
(4.14) $\pi_{1}[1,n{\shortminus}2]s_{n}(s_{0}[2,n{\shortminus}1][1,n{\shortminus}2]s_{n})\cdot\alpha_{n-2}=$ $\ \alpha_{1}.$
Proof.
The lemma follows by direct calculation. ∎
Set
(4.17) $\tilde{\alpha}_{k}=$$\ \left\{\begin{array}[]{r l}\alpha_{k}&\quad\mbox{if }\ 1\leq k\leq n-1,\\ \alpha_{0}+\alpha_{2}+\cdots+\alpha_{n-2}+\alpha_{n}&\quad\mbox{if }\ k=0.\end{array}\right.$
(4.21) $\hat{\alpha}_{k}=$$\ \left\{\begin{array}[]{r l}\alpha_{k}&\quad\mbox{if }\ 1\leq k\leq n-2,\\ \alpha_{n}&\quad\mbox{if }\ k=n-1,\\ \alpha_{0}+\alpha_{2}+\cdots+\alpha_{n-2}+\alpha_{n-1}&\quad\mbox{if }\ k=0.\end{array}\right.$
For $21\leq i\leq i-$ 2, set
$\zeta_{i}=\left\{\begin{array}[]{r l}\big{(}s_{0}[2,n{\shortminus}1][1,n{\shortminus}2]s_{n}\big{)}^{\frac{i}{2}}&\quad\mbox{if}\ i\ \mbox{is even},\\ \pi_{1}[1,n{\shortminus}2]s_{n}\big{(}s_{0}[2,n{\shortminus}1][1,n{\shortminus}2]s_{n}\big{)}^{\frac{i{\shortminus}1}{2}}&\quad\mbox{if}\ i\ \mbox{is odd}.\end{array}\right.$
Lemma 4.2.
For $11\leq k\leq n-$ 1, we have
$\zeta_{i}\cdot\alpha_{k}=\left\{\begin{array}[]{r l}\tilde{\alpha}_{k+i}&\quad\mbox{if}\ i\ \mbox{is even},\\ \hat{\alpha}_{k+i}&\quad\mbox{if}\ i\ \mbox{is odd},\end{array}\right.$
Proof.
Using Lemma 4.1, we observe that the set $\{\tilde{\alpha}_{k}\}_{1\leq k\leq n}$ is closed under the action of $s_{0}[2,n{\shortminus}1][1,n{\shortminus}2]s_{n}$, and that this action sends $\tilde{\alpha}_{k}\mapsto\tilde{\alpha}_{k+1}$ (with indices modulo $n$). Repeating this action $\frac{i}{2}$-times, when $i$ is even, yields the first case of the lemma.
Moreover, it follows that, if $i$ is odd, then $\big{(}s_{0}[2,n{\shortminus}1][1,n{\shortminus}2]s_{n}\big{)}^{\frac{i{\shortminus}1}{2}}$ acts by sending $\tilde{\alpha}_{k}\mapsto\tilde{\alpha}_{k+i-1}$. Using Lemma 4.1 again, we observe that $\pi_{1}[1,n{\shortminus}2]s_{n}$ defines a bijection
$\{\tilde{\alpha}_{k}\}_{1\leq k\leq n}\to\{\hat{\alpha}_{k}\}_{1\leq k\leq n},\qquad\tilde{\alpha}_{k}\mapsto\hat{\alpha}_{k+1}.\qed$
4.1.2. Weights $\omega_{n-1},\omega_{n}$
Set
$\zeta_{i}=\left\{\begin{array}[]{r l}\pi_{n{\shortminus}1}r_{n{\shortminus}2}r_{n{\shortminus}4}\cdots r_{3}s_{n}&\quad\mbox{if }\ i=n{\shortminus}1\ \mbox{is even},\\ \pi_{n{\shortminus}1}s_{n{\shortminus}1}s_{n{\shortminus}2}r_{n{\shortminus}3}r_{n{\shortminus}5}\cdots r_{3}s_{n}&\quad\mbox{if }\ i=n{\shortminus}1\ \mbox{is odd},\\ \pi_{n}r_{n{\shortminus}2}r_{n{\shortminus}4}\cdots r_{4}s_{n}[2,n{\shortminus}1]&\quad\mbox{if }\ i=n\ \mbox{is even},\\ \pi_{n}s_{n{\shortminus}1}s_{n{\shortminus}2}r_{n{\shortminus}3}r_{n{\shortminus}5}\cdots r_{4}s_{n}[2,n{\shortminus}1]&\quad\mbox{if }\ i=n\ \mbox{is odd},\\ \end{array}\right.$
Lemma 4.3.
We have
(4.24) $\zeta_{n{\shortminus}1}\cdot\alpha_{k}=$$\ \left\{\begin{array}[]{r l}\alpha_{k{\shortminus}1}&\quad\mbox{if}\ 2\leq k\leq n{\shortminus}1,\\ \pi_{n{\shortminus}1}[n{\shortminus}1,2]\cdot\alpha_{1}&\quad\mbox{if}\ k=1\ \mbox{is odd},\end{array}\right.$
(4.28) $\zeta_{n}\cdot\alpha_{k}=$$\ \left\{\begin{array}[]{r l}\alpha_{k{\shortminus}1}&\quad\mbox{if}\ 2\leq k\leq n{\shortminus}2,\\ \alpha_{n{\shortminus}2}&\quad\mbox{if}\ k=n,\\ \pi_{n}[n{\shortminus}1,2]\cdot\alpha_{1}&\quad\mbox{if}\ k=1\ \mbox{is odd}.\end{array}\right.$
Proof.
The lemma follows by a direct calculation. ∎
4.2. Type $\mathsf{B}_{n}^{(1)}$
For $1\leq i\leq n$, set
(4.29) $\zeta_{i}=\begin{cases}(s_{0}[2,n][1,n])^{\frac{i}{2}},&\text{if $i$ is even},\\ \pi_{1}[1,n](s_{0}[2,n][1,n])^{\frac{i-1}{2}}&\text{if $i$ is odd}.\end{cases}$
Set
$\tilde{\alpha}_{k}=\begin{cases}\alpha_{k}&1\leq k\leq n-1,\\ \alpha_{0}+2\alpha_{n}+\sum_{j=2}^{n-1}\alpha_{j}&k=0.\end{cases}$
Lemma 4.4.
If $11\leq k\leq n-$ 1 then $\zeta_{i}.\alpha_{k}=\tilde{\alpha}_{i+k}$, where the indices are taken modulo $n$.
Proof.
The proof is similar to that of Lemma 4.3. ∎
4.3. Type $\mathsf{C}_{n}^{(1)}$
Let $\zeta_{i}=(s_{0}s_{1}\cdots s_{n})^{i}$, for $11\leq i\leq n-$ 1, and set
$\tilde{\alpha}_{k}=\begin{cases}\alpha_{k}&1\leq k\leq n-1,\\ \sum_{j=0}^{n}\alpha_{j}&k=0.\end{cases}$
Lemma 4.5.
For $11\leq i,k\leq n-$ 1, we have $\zeta_{i}.\alpha_{k}=\tilde{\alpha}_{i+k}$, where the indices are taken modulo $n$.
Proof.
The proof is similar to that of Lemma 4.3. ∎
5. Good polynomials
The goal of this section is to formulate explicit criteria which will allow us to deduce the compatibility of the new Drinfeld presentations for $\widetilde{\mathbf{U}}$ and $\widetilde{\mathbf{U}}^{\imath}$. From now on assume that $\mathbf{s}=(0,\cdots,0)$ and abbreviate $\widetilde{E}_{i}=-q_{i}^{-2}\mathbb{K}_{i}E_{i}K_{i}^{-1}$.
5.1. Iterated $q$-brackets
As shown in §3, reduced expressions for the fundamental weights in all classical types contain subexpressions almost identical to the formulae encountered in type $\mathsf{A}$. Below we describe the braid group action corresponding to such subexpressions in terms of iterated $q$-brackets.
Define non-commutative polynomials $P_{k}(y_{1},\cdots,y_{k})$ and $P^{\prime}_{k}(y_{1},\cdots,y_{k})$ over ${\mathbb{C}}$ by induction in the following way:
(5.1) $P_{1}(y_{1})=$ $\ y_{1},\qquad$ $P_{k+1}(y_{1},\cdots,y_{k})=$ $\ P_{k}(y_{1},\cdots,y_{k-1},[y_{k},y_{k+1}]_{q}),$
(5.2) $P^{\prime}_{1}(y_{1})=$ $\ y_{1},\qquad$ $P^{\prime}_{k+1}(y_{1},\cdots,y_{k})=$ $\ [P^{\prime}_{k}(y_{1},\cdots,y_{k}),y_{k+1}]_{q}.$
Clearly, we have
(5.3) $P_{k}(y_{1},\cdots,y_{k})=$ $\ P_{l+1}(y_{1},\cdots,y_{l},P_{k-l}(y_{l+1},\cdots,y_{k})),$
(5.4) $P^{\prime}_{k}(y_{1},\cdots,y_{k})=$ $\ P^{\prime}_{k-l+1}(P^{\prime}_{l}(y_{1},\cdots,y_{l}),y_{l+1},\cdots,y_{k}).$
for any $11\leq l\leq k-$ 1.
We say that a tuple $(y_{1},\cdots,y_{k})$ is almost commuting if $y_{m}y_{n}=y_{n}y_{m}$, for $|m-n|>1$.
Lemma 5.1.
The polynomials $P_{k}$ and $P^{\prime}_{k}$ are equal if $(y_{1},\cdots,y_{k})$ is almost commuting. In particular, in that case,
$P_{2}(y_{1},P_{2}(y_{2},y_{3}))=P^{\prime}_{2}(P^{\prime}_{2}(y_{1},y_{2}),y_{3}).$
Proof.
The first statement is proven by induction. By definition, $P_{1}=P^{\prime}_{1}$ and $P_{2}=P^{\prime}_{2}$. Moreover,
(5.5) $P_{k+1}(y_{1},\cdots,y_{k+1})=$ $\ P_{2}(y_{1},P_{k}(y_{2},\cdots,y_{k+1}))$
(5.6) $=$ $\ P_{2}(y_{1},P^{\prime}_{k}(y_{2},\cdots,y_{k+1}))$
(5.7) $=$ $\ P^{\prime}_{k}(P_{2}(y_{1},y_{2}),y_{3},\cdots,y_{k+1})$
(5.8) $=$ $\ P^{\prime}_{k}(P^{\prime}_{2}(y_{1},y_{2}),y_{3},\cdots,y_{k+1})=P^{\prime}_{k+1}(y_{1},\cdots,y_{k+1}),$
where in the second equality we used induction, and in the third equality the fact that the tuple is almost commuting. For the second part, observe that
$P_{2}(y_{1},P_{2}(y_{2},y_{3}))=P_{3}(y_{1},y_{2},y_{3})=P^{\prime}_{3}(y_{1},y_{2},y_{3})=P^{\prime}_{2}(P^{\prime}_{2}(y_{1},y_{2}),y_{3}).\qed$
Lemma 5.2.
If $y_{m}y_{m+1}=y_{m+1}y_{m}$ then
$P_{k}(\cdots,y_{m},y_{m+1},\cdots)=P_{k}(\cdots,y_{m+1},y_{m},\cdots).$
Proof.
It suffices to consider the case $k=3$ with $[y_{1},y_{2}]=0$. Then
(5.9) $P_{3}(y_{1},y_{2},y_{3})=$ $\ y_{1}y_{2}y_{3}-q(y_{1}y_{3}y_{2}+y_{2}y_{3}y_{1})+q^{2}y_{3}y_{2}y_{1}$
(5.10) $=$ $\ y_{2}y_{1}y_{3}-q(y_{1}y_{3}y_{2}+y_{2}y_{3}y_{1})+q^{2}y_{3}y_{1}y_{2}=P_{3}(y_{2},y_{1},y_{3}).\qed$
We say that an ordered tuple $(i_{1},\cdots,i_{k})$ of mutually distinct elements of $\mathbb{I}$ is a chain if $a_{i_{l},i_{l+1}}a_{i_{l+1},i_{l}}=1$ for each $11\leq l\leq k{\shortminus}$ 1. We can identify the Dynkin subdiagram formed by $i_{1},\cdots,i_{k}$ with a Dynkin diagram of type $\mathsf{A}_{k}$. To avoid double subscripts, we relabel $i_{1}=1$, etc. Let us abbreviate
(5.11) $\tau_{l}={[{k{\shortminus}l+1},k]}\cdots[1,l]s_{l}={[{k{\shortminus}l+1},k]}\cdots[2,l+1][1,{l{\shortminus}1}],$
for $1\leq l\leq k$.
Lemma 5.3.
Let $(1,\cdots,k)$ be a chain in $\mathbb{I}$.
  1. (1)
    We have
    $\mathbf{T}_{[k,2]}(B_{1})=P_{k}(B_{1},\cdots,B_{k}).$
    Similarly,
    $T_{[k,2]}(F_{1})=P_{k}(F_{1},\cdots,F_{k}),\qquad T_{[k,2]}(\widetilde{E}_{1})=P_{k}(\widetilde{E}_{1},\cdots,\widetilde{E}_{k}).$
  2. (2)
    For $1\leq l\leq k$, we have
    $\mathbf{T}_{\tau_{l}}(B_{l})=P_{l}(B_{k},\cdots,B_{k-l+2},P_{k-l+1}(B_{1},\cdots,B_{k-l+1})).$
    The same formula holds if $\mathbf{T}_{\tau_{l}}$ is replaced by $T_{\tau_{l}}$, and every $B_{i}$ uniformly replaced by $F_{i}$ or $\widetilde{E}_{i}$.
Proof.
This is proven in [Prz23, Lemmas 9.3–9.4, Proposition 9.6]. ∎
Later, we will also need the polynomials
(5.12) $\mathbf{P}_{i}(a,b)=\sum_{r=0}^{2}q^{r}b^{(2-r)}ab^{(r)}+\mathbb{K}_{i}a,\qquad\widehat{\mathbf{P}}(a,b)=\sum_{r=0}^{2}q^{r}b^{(2-r)}ab^{(r)},$
for $i\in\mathbb{I}$.
5.2. Good polynomials
Let $P(B_{0},\cdots,B_{n})$ be any non-commutative polynomial in the variables $B_{0},\cdots,B_{n}$ with coefficients in ${\mathbb{C}}(\mathbb{K}_{i})_{i\in\mathbb{I}}$. Considering this polynomial as an element of $\widetilde{\mathbf{U}}$, and substituting for an occurence of a letter $B_{j}$ either $\widetilde{E}_{j}$ or $F_{j}$, we can write it as a sum of polynomials which are homogeneous in each of the variables $\widetilde{E}_{j}$ and $F_{j}$:
(5.13) $\eta(P(B_{0},\cdots,B_{n}))=\sum_{\underline{d}}P_{\underline{d}}(\widetilde{E}_{0},F_{0},\cdots,\widetilde{E}_{n},F_{n}),$
where $\underline{d}=(d_{j}^{\pm})_{j\in\mathbb{I}}$, with $d_{j}^{+}=\deg_{\widetilde{E}_{j}}P_{\underline{d}}(\cdot)$ and $d_{j}^{-}=\deg_{F_{j}}P_{\underline{d}}(\cdot)$. [^3] E.g., for $P_{\underline{d}}(\cdot)=\widetilde{E}_{2}^{2}F_{1}$, one has $d_{2}^{+}=2$, $d_{1}^{-}=1$. We call the summands on the RHS subterms, and the corresponding tuple $\underline{d}$ the type of a subterm. We say that that a subterm is mixed if there exist $k,l\in\mathbb{I}$ such that $d_{k}^{+},d_{l}^{-}>0$; otherwise a subterm is called pure.
From now on let us assume that each subterm satisfies $d_{0}^{+}+d_{0}^{-}=1$. Let $P_{+}$ be the sum of all subterms with $d_{0}^{+}=1$, and $P_{-}$ be the sum of all subterms with $d_{0}^{-}=1$. Moreover, let $P_{--}$ be the sum of all subterms which have maximal total degree and $d_{j}^{+}=0$ for all $j\in\mathbb{I}$.
Definition 5.4.
Let $i\in\mathbb{I}_{0}$. A polynomial $P(B_{0},\cdots,B_{n})$ is called:
  1. (1)
    good if every mixed subterm of $P(B_{0},\cdots,B_{n})$ with $d_{0}^{+}>0$ vanishes;
  2. (2)
    $i$-good if it satisfies the following conditions:
    1. (a)
      every subterm $P_{\underline{d}}(\widetilde{E}_{0},F_{0},\cdots,\widetilde{E}_{n},F_{n})$ satisfies [^4] The inequality below should be read as a family of inequalities of the coefficients along each simple root.:
      (5.14) $\sum_{1\leq j\leq n}(d_{j}^{+}+d_{j}^{-})\alpha_{j}\leq\theta-\alpha_{i},\qquad d_{0}^{+}+d_{0}^{-}=1,$
    2. (b)
      $P_{+}=T_{\omega^{\prime}_{i}}(\widetilde{E}_{i})$,  $P_{--}=T_{\omega^{\prime}_{i}}(F_{i})$.
The importance of $i$-good polynomials is conveyed by the following proposition.
Proposition 5.5.
Suppose that $\mathbf{T}_{\omega^{\prime}_{i}}(B_{i})$ can be expressed as an $i$-good polynomial. Then
(5.15) $\eta(\mathbf{T}_{\omega_{i}^{\prime}}(B_{i}))\equiv T_{\omega_{i}^{\prime}}(\eta(B_{i}))\quad\mod\widetilde{\mathbf{U}}_{d_{i}\geq 1,+},$
and the diagram
[…]
commutes modulo $\widetilde{\mathbf{U}}_{\neq i,+}$.
Proof.
Suppose that $\mathbf{T}_{\omega^{\prime}_{i}}(B_{i})=P(B_{0},\cdots,B_{n})$ is an $i$-good polynomial. Definition 5.4 implies that
(5.17) $\eta(P(B_{0},\cdots,B_{n}))=P_{+}+P_{-}=T_{\omega^{\prime}_{i}}(\widetilde{E}_{i})+P_{-}.$
Let us analyse $P_{-}$ further. It can be written as the sum of (i) $P_{--}=T_{\omega^{\prime}_{i}}(F_{i})$, (ii) pure subterms with $\sum_{1\leq j\leq n}d_{j}^{-}\alpha_{j}<\theta-\alpha_{i}$, and (iii) mixed subterms, with $d_{0}^{-}=1$ in each case. It follows from Lemma 2.1 and (5.14) that the degrees of these subterms (in the Drinfeld gradation) are given by
$(i)\ \alpha_{i},\quad(ii)\ \theta-\sum_{1\leq j\leq n}d_{j}^{-}\alpha_{j}>\alpha_{i},\quad(iii)\ \theta+\sum_{1\leq j\leq n}(d_{j}^{+}-d_{j}^{-})\alpha_{j}>\alpha_{i},$
in the respective cases. In particular, the subterms of kinds (ii) and (iii) lie in $\widetilde{\mathbf{U}}_{d_{i}\geq 1,+}$. This completes the proof of (5.15).
Let $\mu$ denote the composition of the eastward and southward arrows, and $\nu$ the composition of the southward and eastward arrows in (5.16). We have $\mu(B_{0})=\nu(B_{0})$ and, by the first part of the proposition, $\mu(B_{-1})\equiv\nu(B_{-1})$ modulo $U_{d_{i}=1,+}$. Since $B_{0}$ and $B_{-1}$ generate $\widetilde{\mathbf{U}}^{\imath}(\widehat{\mathfrak{sl}}_{2})$, and $\mu(B_{0})$ contains terms of degree ±$\pm\alpha_{i}$, it follows that the diagram commutes modulo $\widetilde{\mathbf{U}}_{\neq i,+}$. ∎
6. Auxiliary calculations - braid group actions
In section §7 we will express $\mathbf{T}_{\omega^{\prime}_{i}}(B_{i})$ in terms of certain polynomials. The following calculations will be used to prove that those polynomials are $i$-good.
Lemma 6.1.
If $a_{ji}=-1$, then:
  1. (1)
    $[F_{j},\widetilde{E}_{i}]_{q_{i}}=0$,
  2. (2)
    $T_{i}(F_{j})=[F_{j},F_{i}]_{q_{i}}$ and $T_{i}(\widetilde{E}_{j})=[\widetilde{E}_{j},\widetilde{E}_{i}]_{q_{i}}$.
Proof.
The first part follows from the fact that, by (2.3),
(6.1) $F_{j}E_{i}K_{i}^{-1}-q_{i}E_{i}K_{i}^{-1}F_{j}=q_{i}E_{i}K_{i}^{-1}F_{j}-q_{i}E_{i}K_{i}^{-1}F_{j}=0.$
In the second part, $T_{i}(F_{j})=[F_{j},F_{i}]_{q_{i}}$ follows immediately from the definition of the braid group action, and
(6.2) $T_{i}(\widetilde{E}_{j})=$ $\ -q_{i}^{-2}T_{i}(\mathbb{K}_{j}E_{j}K_{j}^{-1})$
(6.3) $=$ $\ q_{i}^{-3}\mathbb{K}_{i}\mathbb{K}_{j}(E_{j}E_{i}-q_{i}E_{i}E_{j})K_{i}^{-1}K_{j}^{-1}$
(6.4) $=$ $\ q_{i}^{-4}\mathbb{K}_{i}\mathbb{K}_{j}(E_{j}K_{j}^{-1}E_{i}K_{i}^{-1}-q_{i}E_{i}K_{i}^{-1}E_{j}K_{j}^{-1})$
(6.5) $=$ $\ [\widetilde{E}_{j},\widetilde{E}_{i}]_{q_{i}}.\qed$
Lemma 6.2.
If $a_{ij}a_{ji}=2$ then
$\mathbf{T}_{j}\mathbf{T}_{i}(B_{j})=\left\{\begin{array}[]{r l}\mathbb{K}_{i}B_{j}+\sum_{r=0}^{2}(-q)^{r}B_{i}^{(2-r)}B_{j}B_{i}^{(r)}&\quad\mbox{if }\ a_{ji}=-2,\\ \\ {[B_{i},B_{j}]}_{q^{2}}&\quad\mbox{if }\ a_{ji}=-1.\end{array}\right.$
Proof.
Let $a_{ji}=-2$. Then $q_{i}=q$, $q_{j}=q^{2}$ and
${\bf T}_{i}(B_{j})=[2]_{q}^{-1}B_{j}B_{i}^{2}-qB_{i}B_{j}B_{i}+q^{2}[2]_{q}^{-1}B_{i}^{2}B_{j}+\mathbb{K}_{i}B_{j}.$
Moreover, $T_{j}(\mathbb{K}_{j}B_{i})=\mathbb{K}_{j}B_{i}$ and
${\bf T}_{j}([2]_{q}^{-1}B_{j}B_{i}^{2})=$ $\ [2]_{q}^{-1}\mathbb{K}_{j}^{-1}B_{j}B_{i}B_{j}B_{i}B_{j}-q^{2}[2]_{q}^{-1}\mathbb{K}_{j}^{-1}B_{j}B_{i}B_{j}^{2}B_{i}$
$-q^{2}[2]_{q}^{-1}\mathbb{K}_{j}^{-1}B_{j}^{2}B_{i}^{2}B_{j}+q^{4}[2]_{q}^{-1}\mathbb{K}_{j}^{-1}B_{j}^{2}B_{i}B_{j}B_{i},$
${\bf T}_{j}(-qB_{i}B_{j}B_{i})=$ $\ -q\mathbb{K}_{j}^{-1}B_{i}B_{j}^{2}B_{i}B_{j}+q^{3}\mathbb{K}_{j}^{-1}B_{i}B_{j}^{3}B_{i}$
$+q^{3}\mathbb{K}_{j}^{-1}B_{j}B_{i}B_{j}B_{i}B_{j}-q^{5}\mathbb{K}_{j}^{-1}B_{j}B_{i}B_{j}^{2}B_{i},$
${\bf T}_{j}(q^{2}[2]_{q}^{-1}B_{i}^{2}B_{j})=$ $\ q^{2}[2]_{q}^{-1}\mathbb{K}_{j}^{-1}B_{i}B_{j}B_{i}B_{j}^{2}-q^{4}[2]_{q}^{-1}\mathbb{K}_{j}^{-1}B_{i}B_{j}^{2}B_{i}B_{j}$
$-q^{4}[2]_{q}^{-1}\mathbb{K}_{j}^{-1}B_{j}B_{i}^{2}B_{j}^{2}+q^{6}[2]_{q}^{-1}\mathbb{K}_{j}^{-1}B_{j}B_{i}B_{j}B_{i}B_{j}.$
Therefore,
${\bf T}_{j}{\bf T}_{i}(B_{j})=$ $\ q\left(q^{4}+1\right)\mathbb{K}_{j}^{-1}B_{j}B_{i}B_{j}B_{i}B_{j}-q^{2}\left(q^{4}+q^{2}+1\right)[2]_{q}^{-1}\mathbb{K}_{j}^{-1}B_{j}B_{i}B_{j}^{2}B_{i}$
$-q^{2}[2]_{q}^{-1}\mathbb{K}_{j}^{-1}B_{j}^{2}B_{i}^{2}B_{j}+q^{4}[2]_{q}^{-1}\mathbb{K}_{j}^{-1}B_{j}^{2}B_{i}B_{j}B_{i}$
$-\left(q^{4}+q^{2}+1\right)[2]_{q}^{-1}\mathbb{K}_{j}^{-1}B_{i}B_{j}^{2}B_{i}B_{j}+q^{3}\mathbb{K}_{j}^{-1}B_{i}B_{j}^{3}B_{i}$
$q^{2}[2]_{q}^{-1}\mathbb{K}_{j}^{-1}B_{i}B_{j}B_{i}B_{j}^{2}-q^{4}[2]_{q}^{-1}\mathbb{K}_{j}^{-1}B_{j}B_{i}^{2}B_{j}^{2}+\mathbb{K}_{i}B_{j}.$
Repeatedly using the relation
$B_{j}^{2}B_{i}$ $=(q_{j}+q_{j}^{-1})B_{j}B_{i}B_{j}-B_{i}B_{j}^{2}-q_{j}^{-1}\mathbb{K}_{j}B_{i}$
$=(q^{2}+q^{-2})B_{j}B_{i}B_{j}-B_{i}B_{j}^{2}-q^{-2}\mathbb{K}_{j}B_{i},$
we obtain
${\bf T}_{j}{\bf T}_{i}(B_{j})=q^{2}[2]_{q}^{-1}B_{j}B_{i}^{2}+[2]_{q}^{-1}B_{i}^{2}B_{j}-qB_{i}B_{j}B_{i}+\mathbb{K}_{i}B_{j}.$
Next, let $a_{ji}=-1$. Then $q_{i}=q^{2}$ and $q_{j}=q$. We have ${\bf T}_{i}(B_{j})=B_{j}B_{i}-q^{2}B_{i}B_{j}$ and
${\bf T}_{j}{\bf T}_{i}(B_{j})=$ $\ [2]_{q}^{-1}\mathbb{K}_{j}^{-1}B_{j}B_{i}B_{j}^{2}-q\mathbb{K}_{j}^{-1}B_{j}^{2}B_{i}B_{j}+q^{2}[2]_{q}^{-1}\mathbb{K}_{j}^{-1}B_{j}^{3}B_{i}+B_{j}B_{i}$
$-q^{2}[2]_{q}^{-1}\mathbb{K}_{j}^{-1}B_{i}B_{j}^{3}+q^{3}\mathbb{K}_{j}^{-1}B_{j}B_{i}B_{j}^{2}-q^{4}[2]_{q}^{-1}\mathbb{K}_{j}^{-1}B_{j}^{2}B_{i}B_{j}-q^{2}B_{i}B_{j}$
$=$ $\ \left(q^{4}+q^{2}+1\right)[2]_{q}^{-1}\mathbb{K}_{j}^{-1}B_{j}B_{i}B_{j}^{2}-\left(q^{4}+q^{2}+1\right)[2]_{q}^{-1}\mathbb{K}_{j}^{-1}B_{j}^{2}B_{i}B_{j}$
$+q^{2}[2]_{q}^{-1}\mathbb{K}_{j}^{-1}B_{j}^{3}B_{i}+B_{j}B_{i}-q^{2}[2]_{q}^{-1}\mathbb{K}_{j}^{-1}B_{i}B_{j}^{3}-q^{2}B_{i}B_{j}.$
Using the Serre-type relation
$B_{j}^{3}B_{i}=$ $(q^{4}+q^{2}+1)q^{-2}B_{j}^{2}B_{i}B_{j}-\left(q^{4}+q^{2}+1\right)q^{-2}B_{j}B_{i}B_{j}^{2}+B_{i}B_{j}^{3}$
$-[2]_{q}^{2}q^{-1}\mathbb{K}_{j}B_{j}B_{i}+[2]_{q}^{2}q^{-1}\mathbb{K}_{j}B_{i}B_{j},$
it follows that
${\bf T}_{j}{\bf T}_{i}(B_{j})=[B_{i},B_{j}]_{q^{2}},$
completing the proof. ∎
Lemma 6.3.
Let $a_{ij}a_{ji}=2$. If $a_{ji}=-2$ then
(6.6) $\mathbb{K}_{i}\widetilde{E}_{j}+\sum_{r=0}^{2}(-q)^{r}B_{i}^{(2-r)}\widetilde{E}_{j}B_{i}^{(r)}=T_{j}T_{i}(\widetilde{E}_{j})=$ $\ \sum_{r=0}^{2}(-q)^{r}\widetilde{E}_{i}^{(2-r)}\widetilde{E}_{j}\widetilde{E}_{i}^{(r)},$
(6.7) $T_{j}T_{i}(F_{j})=$ $\ \sum_{r=0}^{2}(-q)^{r}F_{i}^{(2-r)}F_{j}F_{i}^{(r)}.$
If $a_{ji}=-1$ then
$[B_{i},\widetilde{E}_{j}]_{q^{2}}=T_{j}T_{i}(\widetilde{E}_{j})=[\widetilde{E}_{i},\widetilde{E}_{j}]_{q^{2}},\qquad T_{j}T_{i}(F_{j})=[F_{i},F_{j}]_{q^{2}}.$
Proof.
We prove the $a_{ji}=-2$ case, leaving the other easier case to the reader. We have $q_{i}=q$ and $q_{j}=q^{2}$. First, observe that
(6.8) $\sum_{r=0}^{2}(-q)^{r}F_{i}^{(2-r)}\widetilde{E}_{j}F_{i}^{(r)}=([2]^{-1}-q^{-1}+q^{-2}[2]^{-1})F_{i}^{2}\widetilde{E}_{j}=0.$
Secondly,
(6.9) $[2]^{-1}(\widetilde{E}_{i}F_{i}+F_{i}\widetilde{E}_{i})\widetilde{E}_{j}-q(\widetilde{E}_{i}\widetilde{E}_{j}F_{i}+F_{i}\widetilde{E}_{j}\widetilde{E}_{i})+q^{2}[2]^{-1}\widetilde{E}_{j}(\widetilde{E}_{i}F_{i}+F_{i}\widetilde{E}_{i})=$
(6.10) $[2]^{-1}(\widetilde{E}_{i}F_{i}+F_{i}\widetilde{E}_{i})\widetilde{E}_{j}-q^{-1}\widetilde{E}_{i}F_{i}\widetilde{E}_{j}-q^{3}\widetilde{E}_{j}F_{i}\widetilde{E}_{i}+q^{2}[2]^{-1}\widetilde{E}_{j}(\widetilde{E}_{i}F_{i}+F_{i}\widetilde{E}_{i})=$
(6.11) $-[2]^{-1}(q^{-2}\widetilde{E}_{i}F_{i}-F_{i}\widetilde{E}_{i})\widetilde{E}_{j}+[2]^{-1}\widetilde{E}_{j}(q^{2}\widetilde{E}_{i}F_{i}-q^{4}F_{i}\widetilde{E}_{i})=$
(6.12) $-q^{-2}[2]^{-1}\mathbb{K}_{i}\big{(}[E_{i},F_{i}]K_{i}^{-1}\widetilde{E}_{j}-q^{4}\widetilde{E}_{j}[E_{i},F_{i}]K_{i}^{-1}\big{)}=$
(6.13) $-q^{-2}(q^{2}-q^{-2})^{-1}\mathbb{K}_{i}\big{(}(1-K_{i}^{-2})\widetilde{E}_{j}-q^{4}\widetilde{E}_{j}(1-K_{i}^{-2})\big{)}=$ $\ -\mathbb{K}_{i}\widetilde{E}_{j}.$
Next, we calculate $T_{j}T_{i}(\widetilde{E}_{j})$. By definition,
$T_{i}(E_{j})=[2]^{-1}E_{i}^{2}E_{j}-q^{-1}E_{i}E_{j}E_{i}+q^{-2}[2]^{-1}E_{j}E_{i}^{2},$
and
$T_{j}T_{i}(E_{j})$ $=-[2]^{-1}E_{j}E_{i}E_{j}E_{i}F_{j}K_{j}+q^{-2}[2]^{-1}E_{j}E_{i}^{2}E_{j}F_{j}K_{j}$
$+q^{-2}[2]^{-1}E_{i}E_{j}^{2}E_{i}F_{j}K_{j}-q^{-4}[2]^{-1}E_{i}E_{j}E_{i}E_{j}F_{j}K_{j}$
$+q^{-1}E_{j}E_{i}F_{j}K_{j}E_{j}E_{i}-q^{-3}E_{j}E_{i}F_{j}K_{j}E_{i}E_{j}-q^{-3}E_{i}E_{j}F_{j}K_{j}E_{j}E_{i}$
$+q^{-5}E_{i}E_{j}F_{j}K_{j}E_{i}E_{j}-q^{-2}[2]^{-1}F_{j}K_{j}E_{j}E_{i}E_{j}E_{i}+q^{-4}[2]^{-1}F_{j}K_{j}E_{j}E_{i}^{2}E_{j}$
$+q^{-4}[2]^{-1}F_{j}K_{j}E_{i}E_{j}^{2}E_{i}-q^{-6}[2]^{-1}F_{j}K_{j}E_{i}E_{j}E_{i}E_{j}.$
Applying relations (2.3)–(2.4), the expression above simplifies to
$T_{j}T_{i}(E_{j})=-q^{-1}E_{i}E_{j}E_{i}+[2]^{-1}E_{j}E_{i}^{2}+q^{-2}[2]^{-1}E_{i}^{2}E_{j}.$
Since $T_{1}T_{2}(\mathbb{K}_{1})=\mathbb{K}_{2}^{2}\mathbb{K}_{1}$ and $T_{1}T_{2}(K_{1})=K_{2}^{2}K_{1}$, we obtain
(6.14) $\ \ \ T_{j}T_{i}(\widetilde{E}_{j})=$ $\ q^{-6}\sum_{r=0}^{2}(-q)^{r}E_{i}^{(2-r)}E_{j}E_{i}^{(r)}K_{i}^{-2}K_{j}^{-1}\mathbb{K}_{i}^{2}\mathbb{K}_{j}=\sum_{r=0}^{2}(-q)^{r}\widetilde{E}_{i}^{(2-r)}\widetilde{E}_{j}\widetilde{E}_{i}^{(r)},$
where the second equality follows from the fact that $E_{i}^{(2-r)}E_{j}E_{i}^{(r)}K_{i}^{-2}K_{j}^{-1}=\linebreak(E_{i}K_{i}^{-1})^{(2-r)}E_{j}K_{j}^{-1}(E_{i}K_{i}^{-1})^{(r)}$ for each $r$.
Now, combining (6.8), (6.13) and (6.14), we obtain the first equality in the statement of the lemma. The second equality is obtained via a similar calculation to that of $T_{j}T_{i}(E_{j})$ above. ∎
7. Weak compatibility
The goal of this section is to prove the following theorem.
Theorem 7.1.
In types $\mathsf{A}_{n}^{(1)},\mathsf{B}_{n}^{(1)},\mathsf{C}_{n}^{(1)},\mathsf{D}_{n}^{(1)}$, for each $1\leq i\leq n$, $\mathbf{T}_{\omega^{\prime}_{i}}(B_{i})$ can be expressed as an $i$-good polynomial. Hence
$\eta(\mathbf{T}_{\omega_{i}^{\prime}}(B_{i}))\equiv T_{\omega_{i}^{\prime}}(\eta(B_{i}))\mod\widetilde{\mathbf{U}}_{d_{i}\geq 1,+}.$
The second statement follows immediately from the first by Proposition 5.5. To prove the first statement, we proceed by a direct, case-by-case method. We first explicitly describe the polynomials expressing $\mathbf{T}_{\omega^{\prime}_{i}}(B_{i})$, using the calculations from §4, and then prove that these polynomials are $i$-good, using the calculations from §6. Type $\mathsf{A}_{n}^{(1)}$ has already been handled in [Prz23, Proposition 9.6].
7.1. Type $\mathsf{D}_{n}^{(1)}$
For convenience, we will sometimes write $B^{\prime}_{l}=B_{l}$ for $22\leq l\leq n{\shortminus}$ 2 and $B^{\prime}_{n{\shortminus}1}=B_{n}$. We use similar notation with $B$ replaced by $F$.
7.1.1. Weights $\omega_{1},\cdots,\omega_{n{\shortminus}2}$
Lemma 7.2.
Let $21\leq i\leq n{\shortminus}$ 2. If $i$ is even, then
(7.1) $\mathbf{T}_{\omega^{\prime}_{i}}(B_{i})=P_{i}(B_{i{\shortminus}1},\cdots,B_{1},P_{n{\shortminus}i}(B_{i+1},\cdots,B_{n{\shortminus}1},P_{n{\shortminus}1}(B_{n},B_{n{\shortminus}2},\cdots,B_{2},B_{0}))).$
If $i$ is odd, then
(7.2) $\mathbf{T}_{\omega^{\prime}_{i}}(B_{i})=P_{i}(B_{i{\shortminus}1},\cdots,B_{1},P_{n{\shortminus}i}(B_{i+1},\cdots,B_{n{\shortminus}2},B_{n},P_{n{\shortminus}1}(B_{n{\shortminus}1},\cdots,B_{2},B_{0}))).$
Proof.
Recall the definitions of $\zeta_{i}$ from §4 and $\tau_{i}$ from (5.11). Since $\omega^{\prime}_{i}=\zeta_{i}\tau_{i}$ and $\ell(\omega^{\prime}_{i})=\ell(\zeta_{i})+\ell(\tau_{i})$, we have $\mathbf{T}_{\omega^{\prime}_{i}}(B_{i})=\mathbf{T}_{\zeta_{i}}\mathbf{T}_{\tau_{i}}(B_{i})$. According to Lemma 5.3 (with $k=n{\shortminus}1$ and $l=i$),
(7.3) $\mathbf{T}_{\tau_{i}}(B_{i})=P_{i}(B_{n{\shortminus}1},\cdots,B_{n{\shortminus}i+1},P_{n{\shortminus}i}(B_{1},\cdots,B_{n{\shortminus}i})).$
Using Lemmas 2.4 and 4.2, we observe that applying $\mathbf{T}_{\zeta_{i}}$ to the RHS of (7.3) merely changes the indices of the variables except the last variable $B_{n{\shortminus}i}$. The precise change of the indices is given by Lemma 4.2, and
$\mathbf{T}_{\zeta_{i}}(B_{n{\shortminus}i})=\left\{\begin{array}[]{r l}P_{n{\shortminus}1}(B_{n},B_{n{\shortminus}2},\cdots,B_{2},B_{0})&\quad\mbox{if}\ i\ \mbox{is even},\\ P_{n{\shortminus}1}(B_{n{\shortminus}1},B_{n{\shortminus}2},\cdots,B_{2},B_{0})&\quad\mbox{if}\ i\ \mbox{is odd},\\ \end{array}\right.$
follows directly from the definition of the braid group action. ∎
Proposition 7.3.
The polynomials (7.1)–(7.2) are $i$-good.
Proof.
We do the even case, the odd one being analogous. First consider the innermost nested polynomial $P_{n{\shortminus}1}(B_{n},B_{n{\shortminus}2},\cdots,B_{2},B_{0})$. Define
$R_{k}=\left\{\begin{array}[]{r l}B_{0}&\quad\mbox{if}\ k=1,\\ P_{k}(B_{k},\cdots,B_{2},B_{0})&\quad\mbox{if}\ 2\leq k\leq n{\shortminus}2,\\ P_{n{\shortminus}1}(B_{n},B_{n{\shortminus}2},\cdots,B_{2},B_{0})&\quad\mbox{if}\ k=n{\shortminus}1.\end{array}\right.$
Let $S_{k}$ be defined analogously, with each $B_{l}$ replaced by $\widetilde{E}_{l}$. We will prove by induction that the polynomials $R_{k}$ are good. Observe that $R_{1}$ is good trivially, and $R_{2}$ is good by part (1) of Lemma 6.1. For $k\geq 2$, we have
$R_{k+1}=P_{3}(B^{\prime}_{k+1},B_{k},R_{k{\shortminus}1}).$
According to the inductive hypothesis, $R_{k{\shortminus}1}$ and $R_{k}$ are good. Therefore, it suffices to show that $P_{2}(F^{\prime}_{k+1},S_{k})=0$. Since $F^{\prime}_{k+1}$ commutes with $S_{k{\shortminus}1}$, Lemma 5.1 implies that
(7.4) $P_{2}(F^{\prime}_{k+1},S_{k})=P^{\prime}_{2}([F^{\prime}_{k+1},\widetilde{E}_{k}]_{q},S_{k{\shortminus}1}),$
which vanishes since $[F^{\prime}_{k+1},\widetilde{E}_{k}]_{q}=0$ by part (1) of Lemma 6.1. We conclude that $R_{k+1}$ is good and, hence, by induction, $R_{n{\shortminus}1}$ is good.
Next, we show that the middle nested polynomial $P_{n{\shortminus}i}(B_{i+1},\cdots,B_{n{\shortminus}1},R_{n{\shortminus}1})$ is good. The proof is similar to the argument above. Set
$R_{k}^{\mathsf{mid}}=\left\{\begin{array}[]{r l}R_{n{\shortminus}1}&\quad\mbox{if}\ k=0,\\ P_{k+1}(B_{n{\shortminus}k},\cdots,B_{n{\shortminus}1},R_{n{\shortminus}1})&\quad\mbox{if}\ 1\leq k\leq n{\shortminus}i{\shortminus}1,\end{array}\right.$
and define $S_{k}^{\mathsf{mid}}$ analogously, with each $B_{l}$ replaced by $\widetilde{E}_{l}$, and $R_{n{\shortminus}1}$ by $S_{n{\shortminus}1}$. We already know that $R_{0}^{\mathsf{mid}}$ is good. The polynomial $R_{1}^{\mathsf{mid}}$ is also good since, by Lemma 5.2,
$P_{2}(F_{n{\shortminus}1},S_{n{\shortminus}1})=P_{2}(\widetilde{E}_{n},P_{2}(F_{n{\shortminus}1},S_{n{\shortminus}2})),$
and $P_{2}(F_{n{\shortminus}1},S_{n{\shortminus}2})=0$ by the same argument as in (7.4).
Next, we claim that $[F_{n{\shortminus}k},S_{k{\shortminus}2}^{\mathsf{mid}}]=0$ for each $12\leq k\leq n{\shortminus}i{\shortminus}$ 1. Since $S_{k{\shortminus}2}^{\mathsf{mid}}=P_{k{\shortminus}1}(\widetilde{E}_{n{\shortminus}k+2},\cdots,\widetilde{E}_{n{\shortminus}1},S_{n{\shortminus}1})$, it suffices to show that $[F_{n{\shortminus}k},S_{n{\shortminus}1}]=0$. This is indeed true because
$[F_{n{\shortminus}k},S_{n{\shortminus}1}]=T_{\zeta_{i}}([F_{n{\shortminus}k{\shortminus}i},\widetilde{E}_{n{\shortminus}i}])=0.$
Therefore,
$P_{2}(F_{n{\shortminus}k},S_{k{\shortminus}1}^{\mathsf{mid}})=P^{\prime}_{2}([F_{n{\shortminus}k},\widetilde{E}_{n{\shortminus}k+1}]_{q},S_{k{\shortminus}2}^{\mathsf{mid}})=0,$
which implies that $R_{k}^{\mathsf{mid}}$ is good. By induction, we conclude that $R_{n{\shortminus}i{\shortminus}1}^{\mathsf{mid}}$ is good.
Finally, we show that the outer polynomial $P_{i}(B_{i{\shortminus}1},\cdots,B_{1},R_{n{\shortminus}i{\shortminus}1}^{\mathsf{mid}})$ is good. Set
$R_{k}^{\mathsf{out}}=\left\{\begin{array}[]{r l}R_{n{\shortminus}i{\shortminus}1}^{\mathsf{mid}}&\quad\mbox{if}\ k=0,\\ P_{k+1}(B_{k},\cdots,B_{1},R_{n{\shortminus}i{\shortminus}1}^{\mathsf{mid}})&\quad\mbox{if}\ 1\leq k\leq i{\shortminus}1,\end{array}\right.$
and define $S_{k}^{\mathsf{out}}$ analogously, with each $B_{l}$ replaced by $\widetilde{E}_{l}$, and $R_{n{\shortminus}i{\shortminus}1}^{\mathsf{mid}}$ by $S_{n{\shortminus}i{\shortminus}1}^{\mathsf{mid}}$. We already know that $R_{0}^{\mathsf{out}}$ is good. The polynomial $R_{1}^{\mathsf{out}}$ is also good since
(7.5) $P_{2}(F_{1},S_{n{\shortminus}i{\shortminus}1}^{\mathsf{mid}})=$ $\ T_{\zeta_{i}}(P_{2}(F_{n{\shortminus}i+1},P_{n{\shortminus}i}(\widetilde{E}_{1},\cdots,\widetilde{E}_{n{\shortminus}i})))$
(7.6) $=$ $\ T_{\zeta_{i}}(P_{n{\shortminus}i}(\widetilde{E}_{1},\cdots,\widetilde{E}_{n{\shortminus}i{\shortminus}1},P_{2}(F_{n{\shortminus}i+1},\widetilde{E}_{n{\shortminus}i}))),$
and $P_{2}(F_{n{\shortminus}i+1},\widetilde{E}_{n{\shortminus}i})=0$ by part (1) of Lemma 6.1. We also claim that $[F_{k},S_{k{\shortminus}2}^{\mathsf{out}}]=0$ for each $12\leq k\leq i{\shortminus}$ 1. Since $S_{k{\shortminus}2}^{\mathsf{out}}=P_{k{\shortminus}1}(\widetilde{E}_{k{\shortminus}2},\cdots,\widetilde{E}_{1},S_{n{\shortminus}i{\shortminus}1}^{\mathsf{mid}})$, it suffices to show that $[F_{k},S_{n{\shortminus}i{\shortminus}1}^{\mathsf{mid}}]=0$. This is indeed true because
$[F_{k},S_{n{\shortminus}i{\shortminus}1}^{\mathsf{mid}}]=T_{\zeta_{i}}([F_{n+k{\shortminus}i},P_{n{\shortminus}i}(\widetilde{E}_{1},\cdots,\widetilde{E}_{n{\shortminus}i})])=0.$
It follows that
$P_{2}(F_{k},S_{k{\shortminus}1}^{\mathsf{out}})=P^{\prime}_{2}([F_{k},\widetilde{E}_{k{\shortminus}1}]_{q},S_{k{\shortminus}2}^{\mathsf{out}})=0,$
which implies that $R_{k}^{\mathsf{out}}$ is good. By induction, we conclude that $R_{i{\shortminus}1}^{\mathsf{out}}$ is good. This proves that (7.1) is a good polynomial.
Next, Lemma 2.4, part (2) of Lemma 5.3, and part (2) of Lemma 6.1 imply that, by uniformly substituting $F_{j}$ or $\widetilde{E}_{j}$ for each $B_{j}$ in the proof of Lemma 7.2, we obtain a proof of the following identities:
(7.7) $T_{\omega^{\prime}_{i}}(F_{i})=$ $\ P_{i}(F_{i{\shortminus}1},\cdots,F_{1},P_{n{\shortminus}i}(F_{i+1},\cdots,F_{n{\shortminus}1},P_{n{\shortminus}1}(F_{n},F_{n{\shortminus}2},\cdots,F_{2},F_{0}))),$
(7.8) $T_{\omega^{\prime}_{i}}(\widetilde{E}_{i})=$ $\ P_{i}(\widetilde{E}_{i{\shortminus}1},\cdots,\widetilde{E}_{1},P_{n{\shortminus}i}(\widetilde{E}_{i+1},\cdots,\widetilde{E}_{n{\shortminus}1},P_{n{\shortminus}1}(\widetilde{E}_{n},\widetilde{E}_{n{\shortminus}2},\cdots,\widetilde{E}_{2},\widetilde{E}_{0}))).$
The fact that (7.1) is a good polynomial, together with the two identities above, yields part (2.b) of the Definition 5.4 of $i$-goodness. Part (2.a) clearly holds as well, which completes the proof. ∎
7.1.2. Weights $\omega_{n-1},\omega_{n}$
We apply similar analysis to weights $\omega_{n-1}$ and $\omega_{n}$.
Proposition 7.4.
We have
(7.9) $\mathbf{T}_{\omega^{\prime}_{n{\shortminus}1}}(B_{n{\shortminus}1})=$ $\ P_{n{\shortminus}1}(B_{n{\shortminus}2},\cdots,B_{1},P_{n{\shortminus}1}(B_{n},B_{n{\shortminus}2},\cdots,B_{2},B_{0})),$
(7.10) $\mathbf{T}_{\omega^{\prime}_{n}}(B_{n})=$ $\ P_{n{\shortminus}1}(B_{n{\shortminus}2},\cdots,B_{1},P_{n{\shortminus}1}(B_{n{\shortminus}1},B_{n{\shortminus}2},\cdots,B_{2},B_{0})).$
The polynomials (7.9)–(7.10) are $i$-good.
Proof.
Let $i\in\{n{\shortminus}1,n\}$. Since $\omega^{\prime}_{i}=\zeta_{i}[1,n{\shortminus}2]$ and $\ell(\omega^{\prime}_{i})=\ell(\zeta_{i})+\ell([1,n{\shortminus}2])$, we have $\mathbf{T}_{\omega^{\prime}_{i}}(B_{i})=\mathbf{T}_{\zeta_{i}}\mathbf{T}_{[1,n{\shortminus}2]}(B_{i})$. According to Lemma 5.3 (with $k=n{\shortminus}1$ and $l=i$),
(7.11) $\mathbf{T}_{[1,n{\shortminus}2]}(B_{i})=P_{n{\shortminus}1}(B^{\prime}_{n{\shortminus}1},B_{n{\shortminus}2},\cdots,B_{1})$
Using Lemma 4.3, we observe that applying $\mathbf{T}_{\zeta_{i}}$ to the RHS of (7.11) merely shifts the indices of the variables by $-1$, except the last variable $B_{1}$, and
$\mathbf{T}_{\zeta_{i}}(B_{1})=\left\{\begin{array}[]{r l}P_{n{\shortminus}1}(B_{n},B_{n{\shortminus}2},\cdots,B_{2},B_{0})&\quad\mbox{if}\ i=n{\shortminus}1,\\ P_{n{\shortminus}1}(B_{n{\shortminus}1},B_{n{\shortminus}2},\cdots,B_{2},B_{0})&\quad\mbox{if}\ i=n.\\ \end{array}\right.$
The proof of $i$-goodness follows the same method as the proof of Proposition 7.3, so we omit it. ∎
7.2. Type $\mathsf{B}_{n}^{(1)}$
In this subsection only we consider polynomials $P_{k}(y_{1},\cdots,y_{k})$ to be defined in the same way as in §5.1 but with $q$ replaced by $q^{2}$. We will also use the polynomials $\mathbf{P}_{i}(a,b)$ from (5.12).
Lemma 7.5.
Let $11\leq i\leq n{\shortminus}$ 1. Then
(7.12) $\mathbf{T}_{\omega^{\prime}_{i}}(B_{i})=P_{i}(B_{i{\shortminus}1},\cdots,B_{1},P_{n{\shortminus}i}(B_{i+1},\cdots,B_{n{\shortminus}1},\mathbf{P}_{n}(P_{n{\shortminus}1}(B_{n{\shortminus}1},\cdots,B_{2},B_{0}),B_{n}))).$
If $i=n$ then
(7.13) $\mathbf{T}_{\omega^{\prime}_{n}}(B_{n})=P_{n}(B_{n-1},\cdots,B_{1},P_{n}(B_{n},\cdots,B_{2},B_{0})).$
Proof.
Let $11\leq i\leq n{\shortminus}$ 1. Arguing in the same way as in the proof of Lemma 7.2, we find that
(7.14) $\mathbf{T}_{\omega^{\prime}_{i}}(B_{i})=\mathbf{T}_{\zeta_{i}}(P_{i}(B_{n{\shortminus}1},\cdots,B_{n{\shortminus}i+1},P_{n{\shortminus}i}(B_{1},\cdots,B_{n{\shortminus}i}))),$
with $\zeta_{i}$ given by (4.29). Lemma 4.4 implies that $\mathbf{T}_{\zeta_{i}}$ shifts the indices of all the variables on the RHS of (7.14) by $i$ modulo $n$, except the last variable $B_{n{\shortminus}i}$. Moreover,
$\mathbf{T}_{\zeta_{i}}(B_{n{\shortminus}i})=\left\{\begin{array}[]{r l}\mathbf{T}_{s_{0}[2,n]}(B_{n{\shortminus}1})&\quad\mbox{if}\ i\ \mbox{is even},\\ \mathbf{T}_{\pi_{1}[1,n]}(B_{n{\shortminus}1})&\quad\mbox{if}\ i\ \mbox{is odd}.\end{array}\right.$
By Lemma 6.2,
(7.15) $\mathbf{T}_{s_{0}[2,n]}(B_{n{\shortminus}1})=$ $\ \mathbf{T}_{s_{0}[2,n-2]}\mathbf{P}_{n}(B_{n{\shortminus}1},B_{n})$
(7.16) $=$ $\ \mathbf{P}_{n}(P_{n{\shortminus}1}(B_{n{\shortminus}1},\cdots,B_{2},B_{0}),B_{n})$
if $i$ is even, with an analogous calculation in the odd case. The proof for $i=n$ is similar to that in type $\mathsf{C}_{n}^{(1)}$, which is done in Lemma 7.7, so we leave it to the reader. ∎
Proposition 7.6.
The polynomials (7.12)–(7.13) are $i$-good.
Proof.
The proof is similar to the proof of Proposition 7.3. Let us explain the differences in the case $11\leq i\leq n-$ 1. Arguing as in the first paragraph of the proof of Proposition 7.3, we deduce that the polynomials $P_{k}(B_{k},\cdots,B_{2},B_{0})$, for $12\leq k\leq n-$ 1, are good. Hence $S_{k}:=P_{k}(B_{k},\cdots,B_{2},\widetilde{E}_{0})=P_{k}(\widetilde{E}_{k},\cdots,\widetilde{E}_{2},\widetilde{E}_{0})$. Then
(7.17) $\mathbf{P}_{n}(S_{n-1},B_{n})=\mathbf{P}_{n}(P_{2}(\widetilde{E}_{n-1},S_{n-2}),B_{n})=P_{2}(\mathbf{P}_{n}(\widetilde{E}_{n-1},B_{n}),S_{n-2}),$
and $\mathbf{P}(\widetilde{E}_{n-1},B_{n})=\widehat{\mathbf{P}}(\widetilde{E}_{n-1},\widetilde{E}_{n})$ by Lemma 6.3. Therefore,
$\mathbf{P}_{n}(S_{n-1},B_{n})=P_{2}(\widehat{\mathbf{P}}(\widetilde{E}_{n-1},\widetilde{E}_{n}),S_{n-2})=\widehat{\mathbf{P}}(S_{n-1},\widetilde{E}_{n}).$
The rest of the argument follows the same pattern as in the proof of 7.3. In particular, we obtain
(7.18) $P_{i}(B_{i{\shortminus}1},\cdots,B_{1},P_{n{\shortminus}i}(B_{i+1},\cdots,B_{n{\shortminus}1},\mathbf{P}_{n}(P_{n{\shortminus}1}(B_{n{\shortminus}1},\cdots,B_{2},\widetilde{E}_{0}),B_{n})))=$
(7.19) $P_{i}(\widetilde{E}_{i{\shortminus}1},\cdots,\widetilde{E}_{1},P_{n{\shortminus}i}(\widetilde{E}_{i+1},\cdots,\widetilde{E}_{n{\shortminus}1},\widehat{\mathbf{P}}(P_{n{\shortminus}1}(\widetilde{E}_{n{\shortminus}1},\cdots,\widetilde{E}_{2},\widetilde{E}_{0}),\widetilde{E}_{n})))=$ $\ T_{\omega^{\prime}_{i}}(\widetilde{E}_{i}),$
(7.20) $\ P_{i}(F_{i{\shortminus}1},\cdots,F_{1},P_{n{\shortminus}i}(F_{i+1},\cdots,F_{n{\shortminus}1},\widehat{\mathbf{P}}(P_{n{\shortminus}1}(F_{n{\shortminus}1},\cdots,F_{2},F_{0}),F_{n})))=$ $\ T_{\omega^{\prime}_{i}}(F_{i}).$
Note that (7.20) is indeed a subterm of (7.12) of maximal total degree, and so $P_{--}=T_{\omega^{\prime}_{i}}(F_{i})$, proving that (7.12) is $i$-good. ∎
7.3. Type $\mathsf{C}_{n}^{(1)}$
Finally, let us consider type $\mathsf{C}_{n}^{(1)}$.
Lemma 7.7.
Let $11\leq i\leq n-$ 1. Then
(7.21) $\mathbf{T}_{\omega^{\prime}_{i}}(B_{i})=P_{i}(B_{i{\shortminus}1},\cdots,B_{1},P_{n{\shortminus}i}(B_{i+1},\cdots,B_{n{\shortminus}1},\mathbb{P}(B_{n},P_{n{\shortminus}1}(B_{n{\shortminus}1},\cdots,B_{2},B_{1},B_{0})))),$
where $\mathbb{P}(a,b)=[a,b]_{q^{2}}$. If $i=n$ then
(7.22) $\mathbf{T}_{\omega^{\prime}_{n}}(B_{n})=\mathbf{P}_{n-1}(\cdots\mathbf{P}_{2}(\mathbf{P}_{1}(B_{0},B_{1}),B_{2}),\cdots,B_{n-1}).$
Proof.
In the first case, the proof is the same as the proof of Lemma 7.5, with the exception that Lemma 4.5 is used instead of Lemma 4.4, and the case of $a_{ji}=-1$ (rather than $a_{ji}=-2$) from Lemma 6.2 is used.
If $i=n$, then one proceeds as follows. One can rewrite (3.5) as
$\omega_{n}=\pi_{n}[n,1]\cdots[n,n-1]s_{n}.$
By Lemma 6.2, $\mathbf{T}_{[n,n-1]}(B_{n})=\mathbf{P}_{n-1}(B_{n},B_{n-1})$. Proceeding by induction, we may assume that
$\mathbf{T}_{[n,k]\cdots[n,n-1]}(B_{n})=\mathbf{P}_{k}(\cdots\mathbf{P}_{n-2}(\mathbf{P}_{n-1}(B_{n},B_{n-1}),B_{n-2}),\cdots,B_{k}).$
By Lemma 2.4, applying $\mathbf{T}_{[n,k-1]}$ shifts the indices of all the $B_{l}$ and $\mathbf{P}_{l}$ by $-1$, except for $B_{n}$, which is transformed into $\mathbf{P}_{n-1}(B_{n},B_{n-1})$. Finally, $\mathbf{T}_{\pi_{n}}$ reverses the indices. ∎
Proposition 7.8.
The polynomials (7.21)–(7.22) are $i$-good.
Proof.
The proof in the case $11\leq i\leq n-$ 1 is similar to those in types $\mathsf{B}_{n}^{(1)}$ and $\mathsf{D}_{n}^{(1)}$, so let us consider the case $i=n$. First, observe that an appropriate modification of the proof of Lemma 7.7 yields
(7.23) ${T}_{[n,k]\cdots[n,n-1]}(\widetilde{E}_{n})=$ $\ \widehat{\mathbf{P}}(\cdots\widehat{\mathbf{P}}(\widehat{\mathbf{P}}(\widetilde{E}_{n},\widetilde{E}_{n-1}),\widetilde{E}_{n-2}),\cdots,\widetilde{E}_{k})=:S_{k},$
(7.24) $T_{\omega^{\prime}_{n}}(\widetilde{E}_{n})=$ $\ \widehat{\mathbf{P}}(\cdots\widehat{\mathbf{P}}(\widehat{\mathbf{P}}(\widetilde{E}_{0},\widetilde{E}_{1}),\widetilde{E}_{2}),\cdots,\widetilde{E}_{n-1}),$
(7.25) $T_{\omega^{\prime}_{n}}(F_{n})=$ $\ \widehat{\mathbf{P}}(\cdots\widehat{\mathbf{P}}(\widehat{\mathbf{P}}(F_{0},F_{1}),F_{2}),\cdots,F_{n-1}),$
for $11\leq k\leq n-$ 1.
We prove by (descending) induction that
(7.26) $R_{k}:=\mathbf{P}_{k}(\cdots\mathbf{P}_{n-2}(\mathbf{P}_{n-1}(\widetilde{E}_{n},B_{n-1}),B_{n-2}),\cdots,B_{k})=S_{k}.$
The base case of $k=n-1$ reduces to Lemma 6.3. Next, suppose that (7.26) holds for $k$. Then
(7.27) $R_{k-1}=\mathbf{P}_{k-1}(R_{k},B_{k-1})=$ $\ \mathbf{P}_{k-1}(S_{k},B_{k-1})$
(7.28) $=$ $\ \mathbb{K}_{k-1}S_{k}+\sum_{r=0}^{2}(-q)^{r}B_{k-1}^{(2-r)}S_{k}B_{k-1}^{(r)}.$
Let us compute the sum above. Observe that $F_{k-1}S_{k}=q^{2}S_{k}F_{k-1}$ (since $F_{k-1}$ commutes with all the letters in $S_{k}$ except for $K_{k}^{-1}$, which appears twice). Hence
(7.29) $\sum_{r=0}^{2}(-q)^{r}F_{k-1}^{(2-r)}S_{k}F_{k-1}^{(r)}=$ $\ [2]^{-1}(q^{4}-(q+q^{-1})q^{3}+q^{2})S_{k}F_{k-1}^{2}=0.$
Moreover, an easy calculation shows that
(7.30) $(F_{k-1}\widetilde{E}_{k-1}+\widetilde{E}_{k-1}F_{k-1})S_{k}\ -$ $\ q(F_{k-1}S_{k}\widetilde{E}_{k-1}+\widetilde{E}_{k-1}S_{k}F_{k-1})$
(7.31) $+$ $\ q^{2}S_{k}(F_{k-1}\widetilde{E}_{k-1}+\widetilde{E}_{k-1}F_{k-1})=-\mathbb{K}_{k-1}S_{k}.$
It follows that $\mathbf{P}_{k-1}(S_{k},B_{k-1})=\widehat{\mathbf{P}}(S_{k},\widetilde{E}_{k-1})=S_{k-1}$, concluding the inductive step.
The induction above yields $R_{1}=S_{1}$. Applying $\mathbf{T}_{\pi_{n}}$, together with (7.24), yields $P_{+}=T_{\omega^{\prime}_{i}}(\widetilde{E}_{i})$, while (7.25) yields $P_{--}=T_{\omega^{\prime}_{i}}(F_{i})$. It follows that (7.22) is $i$-good. ∎
8. Strong compatibility
The goal of this section is to prove a generalization of Theorem 2.5 to the classical types $\mathsf{B}_{n}^{(1)},\mathsf{C}_{n}^{(1)},\mathsf{D}_{n}^{(1)}$.
Theorem 8.1.
Let $\widehat{{\mathfrak{g}}}$ be of type $\mathsf{A}_{n}^{(1)},\mathsf{B}_{n}^{(1)},\mathsf{C}_{n}^{(1)}$ or $\mathsf{D}_{n}^{(1)}$. Then
(8.1) $\eta_{\mathbf{s}}(\boldsymbol{\grave{\Theta}}_{i}(z))\equiv$ $\ \xi_{\mathbf{s}}(\boldsymbol{\grave{\Theta}}_{i}(z))\boldsymbol{\phi}_{i}^{-}(z^{-1})\boldsymbol{\phi}_{i}^{+}(\mathfrak{C}z){\mod\widetilde{\mathbf{U}}_{+}}[\negthinspace[z]\negthinspace],$
(8.2) $\Delta_{\mathbf{s}}(\boldsymbol{\grave{\Theta}}_{i}(z))\equiv$ $\ \eta_{\mathbf{s}}(\boldsymbol{\grave{\Theta}}_{i}(z))\otimes\eta(\boldsymbol{\grave{\Theta}}_{i}(z))\quad\mod{\widetilde{\mathbf{U}}\otimes\widetilde{\mathbf{U}}_{+}}[\negthinspace[z]\negthinspace].$
Type $\mathsf{A}_{n}^{(1)}$ was already handled in [Prz23, Corollary 9.15]. Here we will follow the same methodology as in op. cit. We begin in §8.1 below by recalling results which generalize to our situation without any extra work.
8.1. Generalized factorization
In this subsection let $\widehat{{\mathfrak{g}}}$ be of type $\mathsf{A}_{n}^{(1)},\mathsf{B}_{n}^{(1)},\mathsf{C}_{n}^{(1)}$ or $\mathsf{D}_{n}^{(1)}$. Set $\mathfrak{C}_{i}=\mathfrak{C}^{-1}\mathbb{K}_{i}$.
Lemma 8.2.
For each $i\in\mathbb{I}_{0}$:
$\eta(A_{i,-1})=x^{-}_{i,1}-\mathfrak{C}_{i}x^{+}_{i,-1}K_{i}+Q_{i},\qquad Q_{i}\in\widetilde{\mathbf{U}}_{d_{i}\geq 1,+}.$
Hence
$\eta(A_{i,-1})\equiv\iota_{i}\eta(A_{-1})\quad\mod\widetilde{\mathbf{U}}_{d_{i}\geq 1,+}.$
Moreover,
$\eta(H_{i,1})\equiv\iota_{i}\eta(H_{1})+q_{i}^{2}\mathfrak{C}_{i}^{-1}[Q_{i},F_{i}]_{q_{i}^{-2}}\quad\mod\widetilde{\mathbf{U}}_{d_{i}\geq 2,+}.$
Proof.
See [Prz23, Lemmas 9.8–9.9]. ∎
The statement of Lemma 8.2 should also be seen as the definition of the term $Q_{i}$. As in §5.2, let us write $\mathbf{T}_{\omega^{\prime}_{i}}(B_{i})=P_{+}+P_{-}$. It then follows from the proof of [Prz23, Lemma 9.8] that, explicitly,
(8.3) $Q_{i}=\mathfrak{C}_{i}(P_{-}-P_{--}).$
Proposition 8.3.
Let $r\in{\mathbb{Z}}$. If
(8.4) $[[Q_{i},F_{i}]_{q_{i}^{-2}},x_{i,-r}^{-}]\in\widetilde{\mathbf{U}}_{d_{i}\geq 1,+},$
then
$\eta(A_{i,r})\equiv\iota_{i}\eta(A_{r})\quad\mod\widetilde{\mathbf{U}}_{d_{i}\geq 1,+}.$
Proof.
See [Prz23, Proposition 9.10]. ∎
Lemma 8.4.
If (8.4) is true, then Theorem 8.1 holds.
Proof.
The proof is the same as in [Prz23, Corollary 9.15]. ∎
Therefore, it suffices to prove that condition (8.4) holds in each type.
8.2. Explicit computation
Below we calculate explicitly that (8.4) holds in type $\mathsf{D}_{n}^{(1)}$, for weights $\omega_{1},\cdots,\omega_{n-2}$. The other types and weights can be handled using the same methods.
Given (8.3), an explicit expression for $Q_{i}$ can be obtained from (7.1)–(7.2) [^5] For simplicity, we use formula (7.1). In the odd case, indices $n$ and $n-1$ should be switched.. Substituting either $F_{j}$ or $\widetilde{E}_{j}$ for each occurrence of a variable $B_{j}$, we can write $Q_{i}$ as a sum of polynomials in the variables $F_{j}$ and $\widetilde{E}_{j}$. In the series of lemmas below, we check that condition (8.4) holds for each of these polynomials. Depending on the type of polynomial $P$, we prove one of the following: (i) $P$ vanishes, (ii) $[P,F_{i}]_{q^{-2}}\in\widetilde{\mathbf{U}}_{d_{i}\geq 2,+}$, (iii) $P$ can, modulo $\widetilde{\mathbf{U}}_{d_{i}\geq 2,+}$, be expressed as a polynomials in $x_{j,r}^{+}$ $(j\neq i)$.
Let $\tilde{e}_{i}^{+}=\widetilde{E}_{i}$ and $\tilde{e}_{i}^{-}=F_{i}$. Let us also abbreviate
$R=P_{n{\shortminus}1}(B_{n},B_{n{\shortminus}2},\cdots,B_{2},F_{0}).$
Lemma 8.5.
If $22\leq i\leq n-$ 2, then
$[P_{i}(\widetilde{E}_{i{\shortminus}1},B_{i{\shortminus}2},\cdots,B_{1},P_{n{\shortminus}i}(\widetilde{E}_{i+1},B_{i+2}\cdots,B_{n{\shortminus}1},R)),F_{i}]_{q^{-2}}\in\widetilde{\mathbf{U}}_{d_{i}\geq 2,+}.$
Proof.
We calculate
(8.5) $[P_{i}(\widetilde{E}_{i{\shortminus}1},B_{i{\shortminus}2},\cdots,B_{1},P_{n{\shortminus}i}(\widetilde{E}_{i+1},B_{i+2}\cdots,B_{n{\shortminus}1},R)),F_{i}]_{q^{-2}}=$
(8.6) $[\widetilde{E}_{i-1},P_{i-1}(B_{i-2},\cdots,B_{1},[\widetilde{E}_{i+1},P_{n-i-1}(B_{i+1},\cdots,B_{n-1},[F_{i},R])]].$
For degree reasons, we only need to check that the commutator of $F_{i}$ with
$R^{\prime}=\sum_{\begin{subarray}{c}\varepsilon_{k}\leq\varepsilon_{k+1}\\ \in\{\pm\}\end{subarray}}P_{n{\shortminus}1}(\tilde{e}_{n}^{\varepsilon_{n}},\tilde{e}_{n-2}^{\varepsilon_{n-2}},\cdots,\tilde{e}_{i+1}^{\varepsilon_{i+1}},F_{i},\cdots,F_{2},F_{0})$
vanishes. This reduces to an easily verifiable calculation in the finite quantum group of type $\mathsf{A}_{3}$. ∎
Lemma 8.6.
If $21\leq i\leq n-$ 2, then
(8.7) $P_{i}(B_{i{\shortminus}1},\cdots,B_{1},P_{n{\shortminus}i}(F_{i+1},B_{i+2},\cdots,B_{n{\shortminus}1},P_{n{\shortminus}1}(B_{n},B_{n{\shortminus}2},\cdots,B_{2},F_{0})))$ $=$
(8.8) $\sum_{\varepsilon_{k}\leq\varepsilon_{k+1}}P_{i}(\tilde{e}_{i-1}^{\varepsilon_{i-1}},\cdots,\tilde{e}_{1}^{\varepsilon_{1}},P_{n{\shortminus}i}(F_{i+1},\cdots,F_{n{\shortminus}1},P_{n{\shortminus}1}(F_{n},F_{n{\shortminus}2},\cdots,F_{2},F_{0})))$
(8.9) $+\delta_{i,n-2}\sum_{\varepsilon_{k}\leq\varepsilon_{k+1}}P_{i}(\tilde{e}_{n-3}^{\varepsilon_{i-3}},\cdots,\tilde{e}_{1}^{\varepsilon_{1}},P_{n{\shortminus}i}(\widetilde{E}_{n},P_{n{\shortminus}1}(F_{n-1},F_{n{\shortminus}2},\cdots,F_{2},F_{0})))$
Proof.
By [Prz23, Lemmas 9.3–9.4], the LHS of (8.7) reduces to
$\sum_{\begin{subarray}{c}\varepsilon_{k}\leq\varepsilon_{k+1},\\ \varepsilon^{\prime}_{k}\leq\varepsilon^{\prime}_{k+1}\end{subarray}}P_{i}(\tilde{e}_{i-1}^{\varepsilon^{\prime}_{i-1}},\cdots,\tilde{e}_{1}^{\varepsilon^{\prime}_{1}},P_{n{\shortminus}i}(F_{i+1},F_{i+2}\cdots,F_{n{\shortminus}1},P_{n{\shortminus}1}(\tilde{e}_{n}^{\varepsilon_{n}},\tilde{e}_{n-2}^{\varepsilon_{n-2}},\cdots,\tilde{e}_{2}^{\varepsilon_{2}},F_{0}))).$
The lemma now follows from the following two equalities
(8.10) $[F_{n-1},P_{n{\shortminus}1}(\widetilde{E}_{n},\widetilde{E}_{n{\shortminus}2},B_{n{\shortminus}3},\cdots,B_{2},F_{0})]_{q}=$ $\ 0,$
(8.11) $[F_{n-2},[F_{n-1},P_{n{\shortminus}1}(\widetilde{E}_{n},F_{n{\shortminus}2},B_{n{\shortminus}3},\cdots,B_{2},F_{0})]_{q}]_{q}=$ $\ 0,$
which can be checked directly [^6] Such equalities can easily be verified using, e.g., GAP’s QuaGroup Package [dG03]. (they reduce to calculations in finite quantum groups of type $\mathsf{A}_{3}$ and $\mathsf{D}_{4}$, respectively). ∎
Lemma 8.7.
Let $22\leq i\leq n-$ 2 and
(8.12) $Z=$ $\ \sum_{\varepsilon_{k}\leq\varepsilon_{k+1}}P_{i}(\widetilde{E}_{i-1},\tilde{e}_{i-2}^{\varepsilon_{i-2}},\cdots,\tilde{e}_{1}^{\varepsilon_{1}},P_{n{\shortminus}i}(F_{i+1},\cdots,F_{n{\shortminus}1},P_{n{\shortminus}1}(F_{n},F_{n{\shortminus}2},\cdots,F_{2},F_{0}))).$
The element $[Z,F_{i}]_{q^{-2}}$ can be expressed as a polynomial in $x_{j,r}^{+}$ $(j\neq i)$, with coefficients in ${\mathbb{C}}$. Hence $[[Z,F_{i}]_{q^{-2}},x_{i,s}^{-}]=0$, for all $s\in{\mathbb{Z}}$.
Proof.
The proof is analogous to [Prz23, Lemmas 9.13]. ∎
Let
$Z^{\prime}=\ \sum_{\begin{subarray}{c}\varepsilon_{k}\leq\varepsilon_{k+1},\\ \varepsilon^{\prime}_{k}\geq\varepsilon^{\prime}_{k+1}\end{subarray}}P_{i}(F_{i-1},\cdots,F_{1},P_{n{\shortminus}i}(\widetilde{E}_{i+1},\tilde{e}_{i+2}^{\varepsilon^{\prime}_{i+2}},\cdots,\tilde{e}_{n-1}^{\varepsilon^{\prime}_{n-1}},P_{n{\shortminus}1}(\tilde{e}_{n}^{\varepsilon_{n}},\tilde{e}_{n-2}^{\varepsilon_{n-2}},\cdots,\tilde{e}_{2}^{\varepsilon_{2}},F_{0}))).$
Lemma 8.8.
Let $21\leq i\leq n-$ 2. Then
(8.13) $Z^{\prime}=$ $\ \sum_{\begin{subarray}{c}\varepsilon^{\prime}_{k}\geq\varepsilon^{\prime}_{k+1}\end{subarray}}P_{i}(F_{i-1},\cdots,F_{1},P_{n{\shortminus}i}(\widetilde{E}_{i+1},\tilde{e}_{i+2}^{\varepsilon^{\prime}_{i+2}},\cdots,\tilde{e}_{n-1}^{\varepsilon^{\prime}_{n-1}},P_{n{\shortminus}1}(F_{n},F_{n-2},\cdots,F_{0})))$
(8.14) $+$ $\ \sum_{\begin{subarray}{c}\varepsilon_{k}\leq\varepsilon_{k+1}\end{subarray}}P_{i}(F_{i-1},\cdots,F_{1},P_{n{\shortminus}i}(\widetilde{E}_{i+1},\cdots,\widetilde{E}_{n-1},P_{n{\shortminus}1}(\widetilde{E}_{n},\widetilde{E}_{n-2},\tilde{e}_{n-3}^{\varepsilon_{n-3}},\cdots,\tilde{e}_{2}^{\varepsilon_{2}},F_{0})))$
(8.15) $+$ $\ P_{i}(F_{i-1},\cdots,F_{1},P_{n{\shortminus}i}(\widetilde{E}_{i+1},\cdots,\widetilde{E}_{n-2},\widetilde{E}_{n-1},P_{n{\shortminus}1}(\widetilde{E}_{n},F_{n-2},\cdots,F_{2},F_{0})))$
(8.16) $+$ $\ \delta_{i\neq n-2}P_{i}(F_{i-1},\cdots,F_{1},P_{n{\shortminus}i}(\widetilde{E}_{i+1},\cdots,\widetilde{E}_{n-2},\widetilde{E}_{n},P_{n{\shortminus}1}(F_{n-1},F_{n-2},\cdots,F_{2},F_{0}))).$
Proof.
This follows from (8.10)–(8.11). ∎
Lemma 8.9.
The element $[Z^{\prime},F_{i}]_{q^{-2}}$ can be expressed as a polynomials in $x_{j,r}^{+}$ $(j\neq i)$, with coefficients in ${\mathbb{C}}$, modulo $\widetilde{\mathbf{U}}_{d_{i}\geq 2,+}$. Hence $[[Z^{\prime},F_{i}]_{q^{-2}},x_{i,s}^{-}]=0$ modulo $\widetilde{\mathbf{U}}_{d_{i}\geq 1,+}$, for all $s\in{\mathbb{Z}}$. The same holds if $Z^{\prime}$ is replaced by (8.9).
Proof.
The proof is analogous to [Prz23, Lemmas 9.13]. ∎
9. Application to $q$-characters
We propose a generalization of the notion of $q$-characters to affine quantum symmetric pairs, based on the Lu–Wang Drinfeld-type presentation. We apply Theorem 8.1 to show that our construction is compatible with the usual $q$-character map for quantum affine algebras.
9.1. $q$-characters of quantum affine algebras
Fix $\mathbf{c}=(c_{0},\cdots,c_{n})\in{\mathbb{C}}^{n+1}$ and let $C$ be the image of $\mathfrak{C}$ in $\mathbf{U}^{\imath}_{\mathbf{c}}$. Let $\operatorname{Rep}\mathbf{U}$ be the monoidal category of finite dimensional representations of $\mathbf{U}$. It acts on $\operatorname{Rep}\mathbf{U}^{\imath}_{\mathbf{c}}$, the category of finite dimensional representations of $\mathbf{U}^{\imath}_{\mathbf{c}}$, via the coproduct $\Delta_{\mathbf{c},\mathbf{s}}$ (2.29). It follows from [DK19, Sec. 3.5] (see also [Prz23, Lemmas 2.4]) that any such monoidal action is, up to twisting by a character, equal to the “standard” monoidal action with $\mathbf{s}=(0,\cdots,0)$. Therefore, henceforth we shall assume that we are working with the standard monoidal action.
We will use the same notations for the Grothendieck groups of the categories above. From this point of view, $\operatorname{Rep}\mathbf{U}$ is a ring acting on $\operatorname{Rep}\mathbf{U}^{\imath}_{\mathbf{c}}$. Let $U_{q}(\widetilde{\mathfrak{h}})$ be the subalgebra of $\mathbf{U}$ generated by $h_{i,k}$ ($i\in\mathbb{I}$, $k<0$). Following [FR99, FM01], set
(9.1) $Y_{i,a}=$ $\ K_{\omega_{i}}^{-1}\exp\left(\sum_{k>0}\tilde{h}_{i,-k}a^{k}z^{k}\right)\in{U_{q}(\widetilde{\mathfrak{h}})}[\negthinspace[z]\negthinspace]\qquad(a\in{\mathbb{C}}^{\times}),$
where
(9.2) $\tilde{h}_{i,-k}=\sum_{j\in\mathbb{I}_{0}}\widetilde{C}_{ji}(q^{k})h_{j,-k},$
and $\widetilde{C}(q)$ is the inverse of the $q$-Cartan matrix. Let $\mathcal{Y}={\mathbb{Z}}[Y_{i,a}^{\pm 1}]_{i\in I,a\in{\mathbb{C}}^{\times}}$. By [FR99, Theorem 3], there exists an injective ring homomorphism $\chi_{q}\colon\operatorname{Rep}\mathbf{U}\to\mathcal{Y}\subset{U_{q}(\widetilde{\mathfrak{h}})}[\negthinspace[z]\negthinspace]$, called the $q$-character map, given by
$[V]\mapsto\operatorname{Tr}_{V}\left[\exp\left(-(q-q^{-1})\sum_{i\in\mathbb{I}_{0}}\sum_{k>0}\frac{k}{[k]_{q_{i}}}\pi_{V}(h_{i,k})\otimes\tilde{h}_{i,-k}z^{k}\right)\cdot(\pi_{V}\otimes 1)(T)\right],$
where $T$ is as in [FR99, (3.8)], and $\pi_{V}\colon\mathbf{U}\to\operatorname{End}(V)$ is the representation.
The expression above derives from the Khoroshkin-Tolstoy-Levendorsky-Soibelman-Stukopkin-Damiani (KTLSSD) factorization of the universal $R$-matrix [KT94, LSS93, Dam98]. However, by [FM01, Proposition 2.4], the $q$-character map can equivalently be defined more explicitly in terms of the joint spectrum of the Drinfeld-Cartan operators (i.e., the coefficients of the series $\boldsymbol{\phi}^{\pm}_{i}(z)$). More precisely, there is a one-to-one correspondence between the monomials occurring in $\chi_{q}(V)$ and the common eigenvalues of $\boldsymbol{\phi}^{\pm}_{i}(z)$ on $V$. Let $V=\bigoplus_{\gamma}V_{\gamma}$, with $\gamma=(\gamma^{\pm}_{i,\pm m})_{i\in I,m\in{\mathbb{Z}}_{\geq 0}}$, where
$V_{\gamma}=\{v\in V\mid\exists p\ \forall i\in\mathbb{I}_{0}\ \forall m\in{\mathbb{Z}}_{\geq 0}:(\psi^{\pm}_{i,\pm m}-\gamma^{\pm}_{i,\pm m})^{p}\cdot v=0\},$
be the generalized eigenspace decomposition of $V$. Collect the eigenvalues into generating series $\gamma^{\pm}_{i}(z)=\sum_{m\geq 0}\gamma^{\pm}_{i,\pm m}z^{\pm m}$. By [FM01, Proposition 2.4], the series $\gamma^{\pm}_{i}(z)$ are expansions (at 00 and $\infty$, respectively) of the same rational function of the form
$q_{i}^{\deg Q_{i}-\deg R_{i}}\frac{Q_{i}(q_{i}^{-1}z)R_{i}(q_{i}z)}{Q_{i}(q_{i}z)R_{i}(q_{i}^{-1}z)},$
for some polynomials $Q_{i}(z),R_{i}(z)$ with constant term $1$. Writing
(9.3) $Q_{i}(z)=\prod_{r=1}^{k_{i}}(1-za_{i,r}),\qquad R_{i}(z)=\prod_{s=1}^{l_{i}}(1-zb_{i,s}),$
the $q$-character of $V$ can now be expressed as
$\chi_{q}(V)=\sum_{\gamma}\dim(V_{\gamma})M_{\gamma},\qquad M_{\gamma}=\prod_{i\in\mathbb{I}_{0}}\prod_{r=1}^{k_{i}}Y_{i,a_{i,r}}\prod_{s=1}^{l_{i}}Y_{i,b_{i,s}}^{-1}.$
9.2. Boundary $q$-characters
A two-factor universal $K$-matrix for affine quantum symmetric pairs was defined in [AV24]. However, a KTLSSD-type factorization is not yet available. [^7] In the finite case, a factorization of quasi $K$-matrices was established in [DK19, WZ23]. Therefore, we propose to initiate the study of $q$-characters for affine quantum symmetric pairs based on the Lu–Wang (Drinfeld-type) presentation.
Definition 9.1.
Let
$\mathcal{K}^{0}=\frac{1-q_{i}^{-2}Cz^{2}}{1-Cz^{2}}\exp\left(-(q-q^{-1})\sum_{i\in\mathbb{I}_{0}}\sum_{k>0}\frac{k}{[k]_{q_{i}}}H_{i,k}\otimes\tilde{h}_{i,-k}z^{k}\right)\in{\mathbf{U}^{\imath}_{\mathbf{c}}\otimes U_{q}(\widetilde{\mathfrak{h}})}[\negthinspace[z]\negthinspace].$
We define the boundary $q$-character map to be
$\chi_{q}^{\imath}\colon\operatorname{Rep}\mathbf{U}^{\imath}_{\mathbf{c}}\ \to\ {U_{q}(\widetilde{\mathfrak{h}})}[\negthinspace[z]\negthinspace],\qquad[V]\mapsto\operatorname{Tr}_{V}(\mathcal{K}^{0}\circ(\pi_{V}\otimes 1)).$
Consider ${U_{q}(\widetilde{\mathfrak{h}})}[\negthinspace[z]\negthinspace]$ as a $\mathcal{Y}$-module via the ring homomorphism
(9.4) $\mathcal{Y}\to\mathcal{Y}\hookrightarrow{U_{q}(\widetilde{\mathfrak{h}})}[\negthinspace[z]\negthinspace],\qquad Y_{i,a}\mapsto Y_{i,Ca}Y_{i,a^{-1}}^{-1}.$
Below we apply Theorem 8.1 to show that our boundary $q$-character map is a module homomorphism. First, we need some notation. Given a polynomial $P(z)\in{\mathbb{C}}[z]$ with constant term $1$, let $P^{\dagger}(z)$ be the polynomial with constant term $1$ whose roots are obtained from those of $P(z)$ via the transformation $a\mapsto C^{-1}a^{-1}$; and let $P^{*}(z)$ be the polynomial with constant term $1$ whose roots are the inverses of the roots of $P(z)$.
Proposition 9.2.
Let $W\in\operatorname{Rep}\mathbf{U}$. Then the generalized eigenvalues of $\boldsymbol{\grave{\Theta}}_{i}(z)$ on the restricted representation $\eta_{\mathbf{c}}^{*}(W)$ are of the form:
(9.5) $\gamma_{i}^{\iota}(z)=\frac{\mathbf{Q}_{i}(q_{i}^{-1}z)}{\mathbf{Q}_{i}(q_{i}z)}\frac{\mathbf{Q}_{i}^{\dagger}(q_{i}z)}{\mathbf{Q}_{i}^{\dagger}(q_{i}^{-1}z)},$
where $\mathbf{Q}_{i}(z)$ is a polynomial with constant term $1$. Explicitly,
$\mathbf{Q}_{i}(z)={Q_{i}(Cz)R_{i}^{*}(z)},\quad\mathbf{Q}_{i}^{\dagger}(z)={R_{i}(Cz)Q_{i}^{*}(z)}.$
Proof.
Theorem 8.1 implies that the action of $\boldsymbol{\grave{\Theta}}_{i}(z)$ on $\eta_{\mathbf{c}}^{*}(W)$ is the sum of the action of $\boldsymbol{\phi}_{i}^{-}(z^{-1})\boldsymbol{\phi}_{i}^{+}(Cz)$ and a nilpotent operator. Hence the eigenvalues of $\boldsymbol{\grave{\Theta}}_{i}(z)$ coincide with those of $\boldsymbol{\phi}_{i}^{-}(z^{-1})\boldsymbol{\phi}_{i}^{+}(Cz)$. The result now follows from [FR99, Proposition 1] by the same argument as in the proof of [Prz23, Corollary 5.1]. ∎
Corollary 9.3.
Let $(\mathbf{U},\mathbf{U}^{\imath}_{\mathbf{c}})$ be a split quantum symmetric pair of type $\mathsf{A}_{n}^{(1)},\mathsf{B}_{n}^{(1)},\mathsf{C}_{n}^{(1)}$ or $\mathsf{D}_{n}^{(1)}$. Then the following diagram commutes:
[…]
Proof.
The proof is a modification of the proof of [FM01, Proposition 2.4]. Let $W\in\operatorname{Rep}\mathbf{U}$ and $V\in\operatorname{Rep}\mathbf{U}^{\imath}_{\mathbf{c}}$. We need to calculate $\chi_{q}^{\imath}(V\otimes W)$. Formulae (8.1)–(8.2) in Theorem 8.1, together with Proposition 9.2, imply that the eigenvalues of $\boldsymbol{\grave{\Theta}}_{i}(z)$ on $V\otimes W$ are a product of the eigenvalues of $\boldsymbol{\grave{\Theta}}_{i}(z)$ on $V$, and $\gamma_{i}^{\iota}(z)$. Hence $\operatorname{Tr}_{V\otimes W}(\mathcal{K}^{0}\circ(\pi_{V\otimes W}\otimes 1))=\operatorname{Tr}_{V}(\mathcal{K}^{0}\circ(\pi_{V}\otimes 1))\cdot\operatorname{Tr}_{W}(\mathcal{K}^{0}\circ(\pi_{W}\otimes 1))$, and the eigenvalues of $H_{i,m}$ on $W$ are (up to the overall normalization factor $\frac{1-q_{i}^{-2}Cz^{2}}{1-Cz^{2}}$) of the form
(9.6) $\frac{q_{i}^{m}-q_{i}^{-m}}{n(q-q^{-1})}\left(\sum_{r=1}^{k_{i}}((Ca_{i,r})^{m}-a_{i,r}^{-m})-\sum_{s=1}^{l_{i}}((Cb_{i,s})^{m}-b_{i,s}^{-m})\right),$
keeping the notation from (9.3). Plugging (9.6) into the definition of $\chi_{q}^{\imath}(W)$ and comparing with (9.1), we conclude that $\chi_{q}^{\imath}(W)$ is a linear combination of monomials of the form
$\prod_{i\in\mathbb{I}_{0}}\prod_{r=1}^{k_{i}}Y_{i,Ca_{i,r}}Y_{i,a_{i,r}}^{-1}\prod_{s=1}^{l_{i}}Y_{i,b_{i,s}}Y_{i,Cb_{i,s}}^{-1}.$
Therefore, $\chi_{q}^{\imath}(W)$ is equal to the image of $\chi_{q}(W)$ under (9.4), completing the proof. ∎
Remark 9.4.
In the case of the ordinary $q$-character map $\chi_{q}$, the property of being a ring homomorphism follows almost directly from its definition via the universal $R$-matrix (and the associated properties of transfer matrices). In contrast, in our case, the fact that $\chi_{q}^{\imath}$ intertwines the actions of $\operatorname{Rep}\mathbf{U}$ and $\mathcal{Y}$ is non-trivial, since we are not using a universal $K$-matrix, but working with the Drinfeld-type presentation instead.
Finally, let us consider the rank one example of the $q$-Onsager algebra.
Example 9.5.
We calculate the boundary $q$-characters of restricted evaluation representations, for $(s_{0},s_{1})=(0,0)$. Let $W_{n}(a)$ be the irreducible $n+1$-dimensional representation of $U_{q}(\mathfrak{sl}_{2})$ evaluated at $qa$. Then, by [FR99, (4.3)], its $q$-character is
$\chi_{q}(W_{n}(a))=\sum_{i=0}^{n}M_{i},\qquad M_{i}=\prod_{k=i+1}^{n}Y_{aq^{n-2k+1}}\prod_{k=1}^{i}Y_{aq^{n-2k+3}}^{-1}.$
By Corollary 9.3,
$\chi_{q}^{\imath}(W_{n}(a))=\sum_{i=0}^{n}\mathbf{M}_{i},\quad\mathbf{M}_{i}=\prod_{k=i+1}^{n}Y_{Caq^{n-2k+1}}Y_{a^{-1}q^{-n+2k-1}}^{-1}\prod_{k=1}^{i}Y_{a^{-1}q^{-n+2k-3}}Y_{Caq^{n-2k+3}}^{-1}.$
Writing $\chi_{q}(W_{n}(q^{-2}C^{-1}a^{-1}))=\sum_{i=0}^{n}M^{\prime}_{i}$ and $\chi_{q}^{\imath}(W_{n}(q^{-2}C^{-1}a^{-1}))=\sum_{i=0}^{n}\mathbf{M}^{\prime}_{i}$, it is easy to see that $\mathbf{M}_{i}=\mathbf{M}^{\prime}_{n-i}$, which implies that
$\chi_{q}^{\imath}(W_{n}(a))=\chi_{q}^{\imath}(W_{n}(q^{-2}C^{-1}a^{-1})).$
This is not a coincidence since, by [IT10, Theorem 1.17], $W_{n}(a)$ and $W_{n}(q^{-2}C^{-1}a^{-1})$ are indeed isomorphic as representations of the $q$-Onsager algebra.